### v2

# On the Identifying Power of Generalized Monotonicity

Abstract
In the context of a binary outcome, treatment, and instrument, Balke and Pearl (1993, 1997) establish that the monotonicity condition of Imbens and Angrist (1994) has no identifying power beyond instrument exogeneity for average potential outcomes and average treatment effects in the sense that adding it to instrument exogeneity does not decrease the identified sets for those parameters whenever those restrictions are consistent with the distribution of the observable data. This paper shows that this phenomenon holds in a broader setting with a multi-valued outcome, treatment, and instrument, under an extension of the monotonicity condition that we refer to as generalized monotonicity. We further show that this phenomenon holds for any restriction on treatment response that is stronger than generalized monotonicity provided that these stronger restrictions do not restrict potential outcomes. Importantly, many models of potential treatments previously considered in the literature imply generalized monotonicity, including the types of monotonicity restrictions considered by Kline and Walters (2016), Kirkeboen et al. (2016), and Heckman and Pinto (2018), and the restriction that treatment selection is determined by particular classes of additive random utility models. We show through a series of examples that restrictions on potential treatments can provide identifying power beyond instrument exogeneity for average potential outcomes and average treatment effects when the restrictions imply that the generalized monotonicity condition is violated. In this way, our results shed light on the types of restrictions required for help in identifying average potential outcomes and average treatment effects.
KEYWORDS: Multi-valued Treatments, Average Treatment Effects, Endogeneity, Instrumental Variables.
JEL classification codes: C12, C31, C35, C36
1 Introduction
In their analysis of a setting with a binary outcome, treatment, and instrument, Balke and Pearl (1993, 1997) establish that the monotonicity condition of Imbens and Angrist (1994) has no identifying power beyond instrument exogeneity for average potential outcomes and the average treatment effect (ATE). Here, by no identifying power beyond instrument exogeneity, we mean that adding the monotonicity condition of Imbens and Angrist (1994) to instrument exogeneity does not decrease the identified sets for those parameters whenever those restrictions are consistent with the distribution of the observable data. In this way, their results contrast with the analysis of Imbens and Angrist (1994), who showed that their monotonicity condition and instrument exogeneity permitted identification of the local average treatment effect (LATE). This paper studies the extent to which this phenomenon holds in the broader context of a multi-valued outcome, treatment, and instrument.
Our main result shows that a generalization of the monotonicity condition of Imbens and Angrist (1994) to this richer setting also has no identifying power beyond instrument exogeneity for average potential outcomes and ATEs. We hereafter refer to this condition more succinctly as generalized monotonicity. We show further that this result remains true for any restriction that is in fact stronger than generalized monotonicity provided that these stronger restrictions do not restrict potential outcomes in a sense that we will make precise later. Our analysis thereby accommodates many examples of restrictions on potential treatments that have been considered previously in the literature. In particular, we show that encouragement designs, the types of monotonicity restrictions considered by Kline and Walters (2016), Kirkeboen et al. (2016), and Heckman and Pinto (2018), and certain additive random utility models, including some studied in Lee and Salanié (2023), all satisfy generalized monotonicity. We further establish through a series of examples that models that do not satisfy generalized monotonicity can in fact have identifying power for average potential outcomes and ATEs. In this way, our analysis sheds light on the types of restrictions required for help in identifying average potential outcomes and ATEs.
In establishing our results, we derive the identified sets for average potential outcomes under any such restriction and instrument exogeneity while maintaining the assumption that these restrictions are consistent with the distribution of the observable data. Our derivations reveal that the form of the resulting identified sets parallels the form of those derived by Balke and Pearl (1993, 1997) for a binary outcome, treatment, and instrument. An implication of the form of the identified sets is that average potential outcomes and ATEs are only identified under an identification-at-infinity-type condition when imposing instrument exogeneity and any such restriction.
Our paper differs from the closely related literature that, in the context of a binary outcome, treatment, and instrument, considers the identifying power of the monotonicity condition of Imbens and Angrist (1994) and instrument exogeneity for the distribution (as opposed to the average) of potential outcomes, or considers the identifying power of these conditions when combined with additional restrictions on potential outcomes. In particular, Kamat (2019) shows that the monotonicity condition of Imbens and Angrist (1994) does have identifying power beyond instrument exogeneity for the distribution of potential outcomes. Machado et al. (2019) show that the monotonicity condition of Imbens and Angrist (1994) does have additional identifying power for the ATE beyond instrument exogeneity if one additionally imposes an assumption that requires potential outcomes to vary monotonically with the treatment. Thus, the phenomenon we explore is sensitive to both the choice of parameter and to whether one imposes assumptions on potential outcomes.
The remainder of the paper is organized as follows. Section 2 introduces our formal setup, notation and assumptions, including our generalized monotonicity condition. Our main identification results are presented in Section 3. In Section 4, we provide several examples of restrictions on potential treatments that imply our generalized monotonicity condition, and are thus examples of restrictions that have no identifying power beyond instrument exogeneity for average potential outcomes or ATEs. In contrast, in Section 5, we provide several examples of restrictions on potential treatments that imply that our generalized monotonicity condition does not hold, and further show that some of these restrictions in fact have identifying power beyond instrument exogeneity for average potential outcomes and ATEs. Proofs of all results can be found in the Appendix.
2 Setup and Notation
Denote by $Y\in\mathcal{Y}$ a multi-valued outcome of interest, by $D\in\mathcal{D}$ a multi-valued endogenous regressor, and by $Z\in\mathcal{Z}$ a multi-valued instrumental variable. To rule out degenerate cases, we assume throughout that $2\leq|\mathcal{Y}|<\infty$, $2\leq|\mathcal{D}|<\infty$, and $2\leq|\mathcal{Z}|<\infty$. Further denote by $Y_{d}\in\mathcal{Y}$ the potential outcome if $D=d\in\mathcal{D}$ and by $D_{z}\in\mathcal{D}$ the potential treatment if $Z=z\in\mathcal{Z}$. As usual, we assume that
[…]
Let $P$ denote the distribution of $(Y,D,Z)$ and $Q$ denote the distribution of $((Y_{d}:d\in\mathcal{D}),(D_{z}:z\in\mathcal{Z}),Z)$. We will also denote by $Q|_{Y}$ the corresponding marginal distribution of $(Y_{d}:d\in\mathcal{D})$ under $Q$. For $T$ defined by (1), we have
$(Y,D,Z)=T((Y_{d}:d\in\mathcal{D}),(D_{z}:z\in\mathcal{Z}),Z)~{},$
and therefore $P=QT^{-1}$.
Below we will require that $Q\in\mathbf{Q}$, where $\mathbf{Q}$ is a class of distributions satisfying assumptions that we will specify. Different choices of $\mathbf{Q}$ represent different assumptions that we impose on the distribution of potential outcomes and potential treatments. In this sense, $\mathbf{Q}$ may be viewed as a model for potential outcomes and potential treatments. By analogy with the notation introduced above, we will define $\mathbf{Q}|_{Y}=\{Q|_{Y}:Q\in\mathbf{Q}\}$. We say that $\mathbf{Q}$ does not restrict potential outcomes if and only if $\mathbf{Q}|_{Y}$ equals the entire set of distributions on $\mathcal{Y}^{|\mathcal{D}|}$.
Given $P$ and a model $\mathbf{Q}$, we define the set of $Q\in\mathbf{Q}$ that can rationalize $P$ as
$\mathbf{Q}_{0}(P,\mathbf{Q})=\{Q\in\mathbf{Q}:P=QT^{-1}\}~{}.$
We say $\mathbf{Q}$ is consistent with $P$ if and only if $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$. For every model $\mathbf{Q}$ considered in this paper, every $Q\in\mathbf{Q}$ is assumed to satisfy:
Assumption 2.1 (Instrument Exogeneity).
$((Y_{d}:d\in\mathcal{D}),(D_{z}:z\in\mathcal{Z}))\perp\!\!\!\perp Z$ under $Q$.
Under this assumption, $\mathbf{Q}_{0}(P,\mathbf{Q})$ can be simplified as follows. If $Q$ satisfies Assumption 2.1, then
$p_{yd|z}\equiv P\{Y=y,D=d|Z=z\}=Q\{Y_{d}=y,D_{z}=d|Z=z\}=Q\{Y_{d}=y,D_{z}=d\}~{}.$ (2)
Since the marginal distribution of $Z$ under $P$ and $Q$ are the same, i.e., for all $z\in\mathcal{Z}$,
$P\{Z=z\}=Q\{Z=z\}~{},$
$P=QT^{-1}$ if and only if (2) holds. Thus, if all $Q\in\mathbf{Q}$ satisfies Assumption 2.1, then
$\mathbf{Q}_{0}(P,\mathbf{Q})=\bigg{\{}Q\in\mathbf{Q}:~{}p_{yd|z}=Q\{Y_{d}=y,D_{z}=d\},~{}\forall(y,d,z)\in\mathcal{Y}\times\mathcal{D}\times\mathcal{Z}\bigg{\}}~{}.$
Let $\theta(Q)=(\mathbb{E}_{Q}[Y_{d}]:d\in\mathcal{D})$ denote the vector of average potential outcomes. For fixed $P$ and $\mathbf{Q}$, the identified set for $\theta(Q)$ under $P$ relative to $\mathbf{Q}$ is given by
$\Theta_{0}(P,\mathbf{Q})\equiv\{\theta(Q):Q\in\mathbf{Q}_{0}(P,\mathbf{Q})\}~{}.$
$\Theta_{0}(P,\mathbf{Q})$ is nonempty whenever $\mathbf{Q}_{0}(P,\mathbf{Q})$ is nonempty. By construction, this set is “sharp” in the sense that for any value in the set there exists $Q\in\mathbf{Q}_{0}(P,\mathbf{Q})$ for which $\theta(Q)$ equals the prescribed value. The identified set for $\theta(Q)$ immediately implies that the identified set for any parameter $\lambda=\lambda(\theta)$ is given by $\lambda(\Theta_{0}(P,\mathbf{Q}))$. An important example is $\mathbb{E}_{Q}[Y_{j}]-\mathbb{E}_{Q}[Y_{k}]$, the ATE for treatment $j$ versus treatment $k$.
In the next section, we consider identification in any model of potential treatments that implies the following restriction on all $Q\in\mathbf{Q}$:
Assumption 2.2 (Generalized Monotonicity).
For each $d\in\mathcal{D}$, there exists $z^{\ast}=z^{\ast}(d,Q)\in\mathcal{Z}$ such that
$Q\{D_{z^{\ast}}\neq d,~{}D_{z^{\prime}}=d~{}\mbox{for some}~{}z^{\prime}\neq z^{\ast}\}=0~{}.$ (3)
In what follows, we refer to Assumption 2.2 as generalized monotonicity. It states that under $Q$, for each treatment status $d\in\mathcal{D}$, there exists a value (possibly depending on $d$ and $Q$) of the instrument $z^{\ast}\in\mathcal{Z}$ that maximally encourages all individuals to $d$. Here, by “maximally encourage”, we mean that if an individual does not choose $d$ when $Z=z^{\ast}$, then they never choose $d$ for any other value of $Z$. Equivalently, if an individual chooses $d$ when $Z$ is equal to any value other than $z^{\ast}$, then they have to choose $d$ when $Z=z^{\ast}$. When $\mathcal{D}=\mathcal{Z}=\{0,1\}$, Assumption 2.2 is equivalent to the monotonicity assumption of Imbens and Angrist (1994). In this sense, Assumption 2.2 is a generalization of the monotonicity assumption of Imbens and Angrist (1994) to multiple treatments.
We emphasize that Assumption 2.2 only requires, for each possible value of the treatment, that there exists a value of the instrument that maximally encourages that treatment; it does not require that the value of the instrument is unique. For a given distribution $Q$ and given treatment $d\in\mathcal{D}$, let $\mathcal{Z}^{*}(d,Q)$ denote the set of $z^{\ast}$ that satisfy (3). In this notation, Assumption 2.2 can be restated as $\mathcal{Z}^{*}(d,Q)\neq\emptyset$ for each $d\in\mathcal{D}$. In the statement of Assumption 2.2, $z^{\ast}(d,Q)$ is allowed to change across $Q$. The following lemma shows that $\mathcal{Z}^{*}(d,Q)$ is identified from $P$ and is hence the same for all $Q$ that rationalizes $P$ and satisfies Assumptions 2.1 and 2.2. In what follows, we will therefore write $\mathcal{Z}^{*}(d)$ and $z^{*}(d)$ whenever the given distribution $Q$ rationalizes $P$ and satisfies Assumptions 2.1 and 2.2. This result generalizes the corresponding result in Imbens and Angrist (1994).
Lemma 2.1.
Suppose $Q$ satisfies Assumptions 2.1 and 2.2 and $P=QT^{-1}$. Then, $z\in\mathcal{Z}^{*}(d,Q)$ if and only if
$P\{D=d\mid Z=z\}\geq P\{D=d\mid Z=z^{\prime}\}\text{ for all }z^{\prime}\in\mathcal{Z}~{}.$ (4)
Below we prove the necessity of (4); sufficiency is established in the appendix. Note that, for any $d\in\mathcal{D}$, $z^{\prime}\in\mathcal{Z}$, and any $z\in\mathcal{Z}^{*}(d,Q)$,
$P\{D=d\mid Z=z\}=$ $~{}Q\{D_{z}=d\}$
$=$ $~{}Q\{D_{z}=d,D_{z^{\prime}}=d\}+Q\{D_{z}=d,D_{z^{\prime}}\neq d\}$
$=$ $~{}Q\{D_{z^{\prime}}=d\}+Q\{D_{z}=d,D_{z^{\prime}}\neq d\}$
$\geq$ $~{}Q\{D_{z^{\prime}}=d\}$
$=$ $~{}P\{D=d\mid Z=z^{\prime}\}~{},$
where the first and last equalities exploit Assumption 2.1, and the third equality uses Assumption 2.2.
3 Main Result
In order to describe our main result, we first introduce some further notation. Denote by $\mathbf{Q}_{1}^{\ast}$ the set of all distributions that satisfy Assumption 2.1 and by $\mathbf{Q}_{1,2}^{\ast}$ the set of all distributions that satisfy Assumptions 2.1 and 2.2. In terms of this notation, our main result can be stated as follows:
Theorem 3.1.
Suppose $\mathbf{Q}\subseteq\mathbf{Q}_{1,2}^{\ast}$ and $\mathbf{Q}|_{Y}$ equals the entire set of distributions on $\mathcal{Y}^{|\mathcal{D}|}$. Further suppose $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$. Then, $\Theta_{0}(P,\mathbf{Q})=\Theta_{0}(P,\mathbf{Q}_{1,2}^{\ast})=\Theta_{0}(P,\mathbf{Q}_{1}^{\ast})$.
Theorem 3.1 describes the sense in which restrictions on potential treatments stronger than generalized monotonicity have no identifying power for average potential outcomes and ATEs provided that these stronger restrictions do not restrict potential outcomes. This result is established through Theorems 3.2 and 3.3 below. Theorem 3.2, developed in Section 3.1, characterizes $\Theta_{0}(P,\mathbf{Q})$ for any model $\mathbf{Q}$ that is stronger than Assumptions 2.1 and 2.2, i.e., $\mathbf{Q}\subseteq\mathbf{Q}_{1,2}^{\ast}$, and does not restrict potential outcomes, i.e., for which $\mathbf{Q}|_{Y}$ equals the set of all distributions on $\mathcal{Y}^{|\mathcal{D}|}$. The result shows, in particular, that $\Theta_{0}(P,\mathbf{Q})=\Theta_{0}(P,\mathbf{Q}_{1,2}^{\ast})$ for any such model $\mathbf{Q}$ whenever $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$. Remarkably, this result holds even if the model $\mathbf{Q}$ is strictly more restrictive than Assumptions 2.1 and 2.2 in the sense that $\mathbf{Q}\subsetneqq\mathbf{Q}_{1,2}^{\ast}$. On the other hand, Theorem 3.3, developed in Section 3.2, shows that $\Theta_{0}(P,\mathbf{Q}_{1,2}^{\ast})=\Theta_{0}(P,\mathbf{Q}_{1}^{\ast})$ whenever $\mathbf{Q}_{0}(P,\mathbf{Q}_{1,2}^{\ast})\neq\emptyset$. Together, these results immediately imply Theorem 3.1.
3.1 Identified Sets for $\mathbf{Q}\subseteq\mathbf{Q}_{1,2}^{\ast}$
For $d\in\mathcal{D}$ and $z\in\mathcal{Z}$, define $\beta_{d|z}=\mathbb{E}_{P}[Y\mathbbm{1}\{D=d\}\mid Z=z]$. In addition, define $y^{L}=\min(\mathcal{Y})$ and $y^{U}=\max(\mathcal{Y})$. The following theorem derives the identified set for $\theta(Q)$, relative to any model that assumes instrument exogeneity and generalized monotonicity but does not restrict potential outcomes. Note in particular that the assumptions allow for $\mathbf{Q}\subsetneqq\mathbf{Q}_{1,2}^{\ast}$, in which case the model assumes strictly more than instrument exogeneity and generalized monotonicity.
Theorem 3.2.
Suppose $\mathbf{Q}\subseteq\mathbf{Q}_{1,2}^{\ast}$ and $\mathbf{Q}|_{Y}$ equals the entire set of distributions on $\mathcal{Y}^{|\mathcal{D}|}$. Further suppose $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$. Then,
$\Theta_{0}(P,\mathbf{Q})=\prod_{d\in\mathcal{D}}\left[\beta_{d|z^{*}(d)}+y^{L}(1-\sum_{y\in\mathcal{Y}}p_{yd|z^{*}(d)})),~{}\beta_{d|z^{*}(d)}+y^{U}(1-\sum_{y\in\mathcal{Y}}p_{yd|z^{*}(d)}))\right]~{}.$ (5)
Remark 3.1.
Under the instrument exogeneity and monotonicity assumptions of Imbens and Angrist (1994), Balke and Pearl (1993, 1997) found the same form of the identified set for $\theta(Q)$ as (5) when $\mathcal{Y}=\mathcal{D}=\mathcal{Z}=\{0,1\}$. Theorem 3.2 therefore generalizes the result of Balke and Pearl (1993, 1997) to more than two treatment arms and instrument values, to outcomes taking more than two values, and to show that the same identified set holds when imposing possibly stronger restrictions on potential treatments than generalized monotonicity.
Theorem 3.2 immediately implies the following result on the identified sets for the ATE of treatment $j$ versus $k$:
Corollary 3.1.
Under the assumptions of Theorem 3.2, the identified set for $\mathbb{E}_{Q}[Y_{j}-Y_{k}]$ is given by:
$\biggl{[}(\beta_{j|z^{*}(j)}-\beta_{k|z^{*}(k)})+(y^{L}-y^{U})+y^{U}\sum_{y\in\mathcal{Y}}p_{yk|z^{*}(k)}-y^{L}\sum_{y\in\mathcal{Y}}p_{yj|z^{*}(j)}~{},\\ ~{}(\beta_{j|z^{*}(j)}-\beta_{k|z^{*}(k)})+(y^{U}-y^{L})+y^{L}\sum_{y\in\mathcal{Y}}p_{yk|z^{*}(k)}-y^{U}\sum_{y\in\mathcal{Y}}p_{yj|z^{*}(j)}\biggr{]}~{}.$ (6)
Remark 3.2.
From Corollary 3.1, the width of the identified set for $\mathbb{E}_{Q}[Y_{j}-Y_{k}]$ under the assumptions of Theorem 3.2 is given by
$(y^{U}-y^{L})\left(P\{D\neq j\mid Z=z^{*}(j)\}+P\{D\neq k\mid Z=z^{*}(k)\}\right)~{},$
which, following Lemma 2.1, equals
$(y^{U}-y^{L})\left(\min_{z\in\mathcal{Z}}P\{D\neq j\mid Z=z\}+\min_{z\in\mathcal{Z}}P\{D\neq k\mid Z=z\}\right)~{}.$
Therefore, when imposing generalized monotonicity, as well as when imposing any restriction implying generalized monotonity, the ATE of $j$ versus $k$ is only identified “at infinity” (Heckman, 1990; Andrews and Schafgans, 1998) in the sense that identification requires
$\min_{z\in\mathcal{Z}}P\{D\neq j\mid Z=z\}=\min_{z\in\mathcal{Z}}P\{D\neq k\mid Z=z\}=0~{}.$
In other words, identification of the the ATE of $j$ versus $k$ under generalized monotonicity or under any restriction implying generalized monotonicity requires that there is some value of the instrument such that everyone takes treatment $j$ at that value of the instrument, and some value of the instrument such that everyone takes treatment $k$ at that value of the instrument.
3.2 Identifying Power of Generalized Monotonicity
Theorem 3.2 above establishes that, for possibly multi-valued $Y$, $D$, and $Z$, and any model $\mathbf{Q}\subseteq\mathbf{Q}_{1,2}^{\ast}$ such that $\mathbf{Q}|_{Y}$ equals the entire set of distributions on $\mathcal{Y}^{|\mathcal{D}|}$, if $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$, then $\Theta_{0}(P,\mathbf{Q})$ equals (5). We now show that $\Theta_{0}(P,\mathbf{Q}_{1,2}^{\ast})=\Theta_{0}(P,\mathbf{Q}_{1}^{\ast})$ as long as $\mathbf{Q}_{0}(P,\mathbf{Q}_{1,2}^{\ast})\neq\emptyset$, so that the identified set for $\theta(Q)$ assuming instrument exogeneity and generalized monotonicity coincides with the identified set assuming instrument exogeneity alone, as long as both assumptions are consistent with the distribution of the observed data. Our result therefore generalizes Balke and Pearl (1993, 1997), which study the case of binary $Y$, $Z$, and $D$.
In order to do so, we consider a mean independence assumption even weaker than Assumption 2.1, and show the identified set under this even weaker assumption is also (5). A sandwich argument will then lead to our desired result. In particular, we first establish that the identified set for $\theta(Q)$ in (5) coincides with the identified set under the weaker mean independence assumption considered by Robins (1989) and Manski (1990):
Assumption 3.1.
$\mathbb{E}_{Q}[Y_{d}\mid Z=z]=\mathbb{E}_{Q}[Y_{d}]$ for all $d\in\mathcal{D}$ and $z\in\mathcal{Z}$.
Note Assumption 3.1 is weaker than instrument exogeneity in Assumption 2.1, and does not imply (2). Let $\mathbf{Q}_{3}^{\ast}$ denote the set of all $Q$ that satisfies Assumption 3.1. Following Robins (1989) and Manski (1990), the following lemma derives the identified set for $\theta(Q)$ under mean independence:
Lemma 3.1.
Suppose $\mathbf{Q}_{0}(P,\mathbf{Q}_{3}^{\ast})\neq\emptyset$. Then,
$\Theta_{0}(P,\mathbf{Q}_{3}^{\ast})=\prod_{d\in\mathcal{D}}\left[\max_{z\in\mathcal{Z}}\{\beta_{d|z}+y^{L}(1-\sum_{y\in\mathcal{Y}}p_{yd\mid z})\},~{}\min_{z\in\mathcal{Z}}\{\beta_{d|z}+y^{U}(1-\sum_{y\in\mathcal{Y}}p_{yd\mid z})\}\right].$ (7)
The following lemma, which relies on the observation in Lemma 2.1, establishes the equivalence between the identified sets in (5) and (7) when $\mathbf{Q}_{0}(P,\mathbf{Q}_{1,2}^{\ast})\neq\emptyset$.
Lemma 3.2.
Suppose $\mathbf{Q}_{0}(P,\mathbf{Q}_{1,2}^{\ast})\neq\emptyset$. Then, the sets in (5) and (7) coincide.
Using Lemma 3.2, we are able to establish our desired result, which asserts that, maintaining Assumption 2.1, additionally imposing Assumption 2.2 either causes the identified set for $\theta(Q)$ to become empty (if those assumptions are not consistent with $P$) or leaves the identified set for $\theta(Q)$ unchanged (if those assumptions are consistent with $P$).
Theorem 3.3.
Suppose $\mathbf{Q}_{0}(P,\mathbf{Q}_{1,2}^{\ast})\neq\emptyset$. Then, $\Theta_{0}(P,\mathbf{Q}_{1,2}^{\ast})=\Theta_{0}(P,\mathbf{Q}_{1}^{\ast})$.
4 Examples of Models That Satisfy Assumption 2.2
We now consider some restrictions on potential treatments that have been considered previously in the literature. In each case, we show $\mathbf{Q}\subseteq\mathbf{Q}_{1,2}^{\ast}$; in particular, these restrictions satisfy generalized monotonicity. We emphasize that frequently $\mathbf{Q}\subsetneqq\mathbf{Q}_{1,2}^{\ast}$. In what follows, it is implicitly understood that $\mathbf{Q}|_{Y}$ equals the entire set of distributions on $\mathcal{Y}^{|\mathcal{D}|}$, so that the model imposes no restriction on potential outcomes. Thus, in each example Theorem 3.1 applies and such restrictions do not provide any identifying power for average potential outcomes or ATEs.
Example 4.1.
Consider a multi-arm randomized controlled trial (RCT) with noncompliance, where $Z=d$ denotes random assignment to treatment $d$, $D_{d}=d$ denotes that the subject would comply with assignment if assigned to treatment $d$, and $|\mathcal{D}|=|\mathcal{Z}|$. More generally, not necessarily in the context of an RCT, one can interpret $Z=d$ as encouragement to treatment $d$ and interpret $D_{d}=d$ as the subject would take treatment $d$ if encouraged to do so. In this example, $Q$ satisfies Assumption 2.1 because $Z$ is randomly assigned. We may generalize the “no-defier” restriction of of Angrist et al. (1996) as follows:
$Q\{D_{d}\neq d,~{}D_{d^{\prime}}=d~{}\mbox{for some}~{}d^{\prime}\neq d\}=0~{},$
i.e., there is zero probability that a subject would not take treatment $d$ if assigned to (encouraged to take) $d$ but would take $d$ if assigned (encouraged) to some other treatment $d^{\prime}\neq d$. If $Q$ satisfies this generalized no-defier restriction, then Assumption 2.2 holds with $z^{*}(d)=d$ for all $d$. This no-defier restriction in particular holds in the context of an RCT with “one-sided non-compliance,” where we assume
$Q\{D_{z}\in\{0,z\}\}=1~{},$
for all $z\in\mathcal{Z}$. Here, non-compliance is one-sided because one can fall back to the control group if assigned to $d$ but cannot choose $d$ if assigned to the control group.
Example 4.2.
Cheng and Small (2006) consider an RCT with noncompliance where $\mathcal{D}=\mathcal{Z}=\{0,1,2\}$. Assumption 2.1 continues to hold because $Z$ is randomly assigned. In such a setting, they develop bounds on average effects within subgroups defined by potential treatments which, following the terminology of Frangakis and Rubin (2002), they call “principal strata.” While Bai et al. (2024) derives the identified sets for those parameters given their assumptions, we now use our analysis to consider instead identification of average potential outcomes and ATEs given their assumptions. Their “Monotonicity I” assumption is equivalent to one-sided noncompliance in the preceding example. Their “Monotonicity II” assumption states that subjects who would comply with assignment to treatment $2$ would also comply with assignment to treatment $1$, so that
$Q\{D_{1}=1\mid D_{2}=2\}=1~{}.$
They argue that such an assumption is plausible in a medical context when treatment $1$ has fewer side effects than treatment $2$, and in their application to treatments for alcohol dependence, in which complying with treatment $1$ (compliance enhancement therapy) requires less effort by subjects than complying with treatment $2$ (cognitive behavioural therapy). Because their Monotonicity I restriction implies our Assumption 2.2, so does imposing both their Monotonicity I and II restrictions.
Example 4.3.
Kline and Walters (2016) considers an RCT with a “close substitute” to study the effects of preschooling on educational outcomes. In their setting, $D\in\mathcal{D}=\{0,1,2\}$, where $D=0$ denotes home care (no preschool), $D=2$ denotes a preschool program called Head Start, and $D=1$ denotes preschools other than Head Start, namely the close substitute. Let $Z\in\mathcal{Z}=\{0,1\}$ denote an indicator variable for an offer to attend Head Start. Assumption 2.1 holds because $Z$ is randomly assigned. Kline and Walters (2016) impose the restriction that
$Q\{D_{1}=2\mid D_{0}\neq D_{1}\}=1~{}.$ (8)
The condition in (8) states that if the choice of a family changes upon receiving a Head Start offer, then they must choose Head Start when receiving the offer. In other words, it cannot be the case that upon receiving a Head Start offer, a family switches from no preschool to preschools other than Head Start, or the other way around. Assumption 2.2 then holds with $z^{*}(0)=z^{*}(1)=0$ and $z^{*}(2)=1$. To see this, note that (8) implies $Q\{D_{0}\neq D_{1},D_{1}\neq 2\}=0$ and thus
$Q\{D_{0}\neq 0,D_{1}=0\}$ $=Q\{D_{0}\neq D_{1},D_{1}=0\}=0~{},$
$Q\{D_{0}\neq 1,D_{1}=1\}$ $=Q\{D_{0}\neq D_{1},D_{1}=1\}=0~{},$
$Q\{D_{1}\neq 2,D_{0}=2\}$ $\leq Q\{D_{0}\neq D_{1},D_{1}\neq 2\}=0~{}.$
Note in this example Assumption 2.2 still holds although $|\mathcal{Z}|<|\mathcal{D}|$.
Example 4.4.
Suppose $Q$ satisfies Assumption 2.1. Heckman and Pinto (2018) define “unordered monotonicity” as the assumption that, for any $d\in\mathcal{D}$, and any $z,z^{\prime}\in\mathcal{Z}$,
$Q\{\mathbbm{1}\{D_{z}=d\}\geq\mathbbm{1}\{D_{z^{\prime}}=d\}\}=1\text{ or }Q\{\mathbbm{1}\{D_{z}=d\}\leq\mathbbm{1}\{D_{z^{\prime}}=d\}\}=1~{}.$ (9)
Assumption 2.2 holds for any $Q$ that satisfies (9). To see this, note that unordered monotonicity immediately implies that, for each $d\in\mathcal{D}$, $\mathbbm{1}\{D_{z}=d\}$ is weakly ordered across all $z\in\mathcal{Z}$ (w.p.1), and Assumption 2.2 is immediately seen to hold by taking $z^{\ast}(d)$ to be the maximum with respect to this order. Note, however, that $\mathcal{Z}^{*}(d)$ may not be a singleton unless some inequalities in (9) are strict.
Remark 4.1.
Although unordered monotonicity implies Assumption 2.2, the converse is generally false. For example, suppose $\mathcal{Z}=\{0,1,2,3\}$ and $\mathcal{D}=\{0,1\}$. Suppose $\mathbbm{1}\{D_{3}=1\}\geq\mathbbm{1}\{D_{z}=1\}$ w.p.1 under $Q$ for $z\neq 3$ and $\mathbbm{1}\{D_{0}=0\}\geq\mathbbm{1}\{D_{z}=0\}$ w.p.1 under $Q$ for $z\neq 0$, but $Q\{D_{1}=1,D_{2}=0\}$ and $Q\{D_{1}=0,D_{2}=1\}$ are both strictly positive. Then Assumption 2.2 holds with $z^{*}(0)=0$ and $z^{*}(1)=3$, but unordered monotonicity fails. In particular, $\mathbbm{1}\{D_{1}=1\}$ and $\mathbbm{1}\{D_{2}=1\}$ are not ordered, thus violating (9). We thus conclude that if $\mathbf{Q}$ is defined as the set of distributions that satisfy instrument exogeneity and unordered monotonicity, then $\mathbf{Q}\subsetneqq\mathbf{Q}_{1,2}^{\ast}$.
Example 4.5.
Suppose under $Q$, $(D_{z}:z\in\mathcal{Z})$ is determined by
$D_{z}=\operatorname*{argmax}_{d\in\mathcal{D}}~{}(g(z,d)+U_{d})~{},$ (10)
for $g:\mathcal{Z}\times\mathcal{D}\to\Re$ where $\Re$ is the set of real numbers and a random vector $(U_{d}:d\in\mathcal{D})$, whose distribution is absolutely continuous with respect to the Lebesgue measure on $\Re^{|\mathcal{D}|}$ and $Z\perp\!\!\!\perp((U_{d}:d\in\mathcal{D}),(Y_{d}:d\in\mathcal{D}))$. Hence, $Q$ satisfies Assumption 2.1 by construction. Let $\mathbf{Q}$ denote the set of distributions that are consistent with $(D_{z}:z\in\mathcal{Z})$ being determined by (10) for some $g$ and $(U_{d}:d\in\mathcal{D})$ satisfying these requirements. The model $\mathbf{Q}$ is called an additive random utility model (ARUM). A sufficient condition for $Q\in\mathbf{Q}$ to satisfy Assumption 2.2 is that for each $d\in\mathcal{D}$ there exists $z^{*}(d)\in\mathcal{Z}$ such that
$g(z^{*}(d),d)-g(z^{*}(d),d^{\prime})>g(z,d)-g(z,d^{\prime})\text{ for all }d^{\prime}\neq d\text{ and }z\neq z^{*}(d)~{}.$ (11)
We refer to the requirement in (11) as uniform targeting of treatment $d$. The terminology is intended to reflect that there is a value of the instrument that maximizes the gains (in terms of $g$) of choosing treatment $d$ versus any other treatment $d^{\prime}$ uniformly across these other possible values of the treatment. In this sense, that value of the instrument targets treatment $d$ uniformly. We now argue by contradiction that (11) implies (3); hence, if (11) holds for all $d\in\mathcal{D}$, then $Q$ satisfies Assumption 2.2. To this end, suppose that, with positive probability, $D_{z^{*}(d)}=d^{\prime}\neq d$ but $D_{z^{\prime}}=d$ for $z^{\prime}\neq z^{*}(d)$. Then,
$g(z^{*}(d),d^{\prime})+U_{d^{\prime}}$ $\geq g(z^{*}(d),d)+U_{d}~{},$
$g(z^{\prime},d)+U_{d}$ $\geq g(z^{\prime},d^{\prime})+U_{d^{\prime}}~{}.$
These two inequalities imply
$g(z^{*}(d),d)-g(z^{*}(d),d^{\prime})\leq g(z^{\prime},d)-g(z^{\prime},d^{\prime})~{},$
which violates (11). A particular example when $|\mathcal{Z}|\geq|\mathcal{D}|$ that satisfies (11) with $z^{*}(d)=d$ after a suitable relabelling is
$g(z,d)=\alpha_{d}+\beta_{d}\mathbbm{1}\{z=d\}~{},$ (12)
with $\beta_{d}>0$, so that $Z=d$ strictly increases the latent value of treatment $d$ while leaving the values of the remaining options unchanged.
Example 4.6.
Consider the ARUM of Example 4.5 when $|\mathcal{D}|=2$, and let $\mathbf{Q}$ denote the set of distributions defined in that example. Then, Assumption 2.2 holds for all $Q\in\mathbf{Q}$. To see this, consider $Q\in\mathbf{Q}$. Label $\mathcal{D}=\{0,1\}$, and let $g_{10}(z)=g(z,1)-g(z,0)$ and $U_{10}=U_{1}-U_{0}$. The assumptions of Example 4.5 on $(U_{1},U_{0})$ imply that the distribution of $U_{10}$ is absolutely continuous with respect to Lebesgue measure and that $U_{10}\perp\!\!\!\perp Z$. Ignoring ties that occur with probability zero, (10) can be rewritten as
$D_{z}=\mathbbm{1}\{g_{10}(z)+U_{10}\geq 0\}~{}.$ (13)
Let $\overline{\mathcal{Z}}=\operatorname*{argmax}_{z\in\mathcal{Z}}\{g_{10}(z)\}$, and let $\underline{\mathcal{Z}}=\operatorname*{argmin}_{z\in\mathcal{Z}}\{g_{10}(z)\}$. Then, $Q$ satisfies Assumption 2.2 with $\mathcal{Z}^{*}(1)=\overline{\mathcal{Z}}$ and $\mathcal{Z}^{*}(0)=\underline{\mathcal{Z}}$. To contrast with Example 4.5, note that (11) holds if and only if $\overline{\mathcal{Z}}$ and $\underline{\mathcal{Z}}$ are both singletons.
Remark 4.2.
The ARUM for a binary treatment of (13) is equivalent to the Heckman-Vytlacil nonparametric selection model for a binary treatment considered, e.g., in Heckman and Vytlacil (1999, 2005), and shown by Vytlacil (2002) to be equivalent to the monotonicity and exogeneity assumptions of Imbens and Angrist (1994). Heckman and Vytlacil (2001) show that the Heckman-Vytlacil nonparametric selection model for a binary treatment results in an identified set for the ATE in (6) and that the model has no identifying power beyond instrument exogeneity for the ATE. Example 4.6 therefore shows that the results of this paper nests the results of Heckman and Vytlacil (2001); Example 4.5, on the other hand, extends their results to nonparametric selection models for a multi-valued treatment. For a partial identification analysis of a class of parameters that includes the ATE under a nonparametric selection model for a binary treatment and sometimes imposing additional restrictions, see, e.g., Mogstad et al. (2018), Han and Yang (2024), and Marx (2024).
Example 4.7.
Lee and Salanié (2023) also consider the ARUM defined by (10) without imposing the uniform targeting of (11) for each treatment. Instead, they impose an assumption that they refer to as “strict one-to-one targeting”, in which the set of treatments can be partitioned into a set of treatments $\mathcal{D}^{\dagger}$ that are “targeted” and a set of treatments $\mathcal{D}\setminus\mathcal{D}^{\dagger}$ that are “not targeted” such that
  1. 1.
    For $d\in\mathcal{D}\setminus\mathcal{D}^{\dagger}$, $g(z,d)$ is the same for all $z\in\mathcal{Z}$;
  2. 2.
    For $d\in\mathcal{D}^{\dagger}$, there exists $z^{\dagger}(d)$ such that $g(z^{\dagger}(d),d)>g(z^{\prime},d)$ for all $z^{\prime}\neq z^{\dagger}(d)$ and such that $g(z^{\prime},d)$ takes the same value for all $z^{\prime}\neq z^{\dagger}(d)$; additionally, $z^{\dagger}(d)\neq z^{\dagger}(d^{\prime})$ for $d,d^{\prime}\in\mathcal{D}^{\dagger}$, $d\neq d^{\prime}$.
The terminology “one-to-one” stems from the second requirement above. They further impose that there exists a treatment that is known to be non-targeted. This class of ARUMs is equivalent to imposing (12) for targeted treatments, imposing $g(z,d)=\alpha_{d}$ for non-targeted treatments, and imposing that there is at least one non-targeted treatment. In such a setting, Lee and Salanié (2023) analyze the identification of a particular class of average effects within subgroups defined by potential treatments. We now use our analysis to consider instead the identification of average potential outcomes and ATEs given their assumptions. Suppose strict one-to-one targeting holds, and additionally suppose that there is at least one targeted treatment. In Appendix B.1, we argue that Assumption 2.2 holds when $|\mathcal{Z}|>|\mathcal{D}^{\dagger}|$, so that there are more values of the instrument than there are targeted treatments. We further argue that (11) does not hold for some treatments unless $|\mathcal{D}|=|\mathcal{Z}|=2$. Such models therefore provide another class of ARUMs, distinct from the one with uniform targeting described in Example 4.5, for which Assumption 2.2 holds. In Example 5.3 below, we show, however, that Assumption 2.2 does not hold when $|\mathcal{D}|\geq 3$, $|\mathcal{Z}|=|\mathcal{D}^{\dagger}|$, and the support of $(U_{d}:d\in\mathcal{D})$ is $\Re^{|\mathcal{D}|}$.
Example 4.8.
Kirkeboen et al. (2016) study the effects of fields of study on earnings. In their setting, $\mathcal{D}=\{0,1,2\}$ represent three fields of study, ordered by their (soft) admission cutoffs from the lowest to the highest. The instrument is $Z\in\{0,1,2\}$, with $Z=1$ when the student crosses the (soft) admission cutoff for field 1, $Z=2$ when the student crosses the (soft) admission cutoff for field 2, and $Z=0$ otherwise. The authors assume that $Z$ is exogenous in the sense that $Q$ satisfies Assumption 2.1 and impose the following monotonicity conditions:
$Q\{D_{1}=1|D_{0}=1\}$ $=1~{},$ (14)
$Q\{D_{2}=2|D_{0}=2\}$ $=1~{}.$ (15)
The conditions in (14)–(15) require that crossing the cutoff for field 1 or 2 weakly encourages them towards that field. They further impose the following “irrelevance” conditions:
$Q\{\mathbbm{1}\{D_{1}=2\}$ $=\mathbbm{1}\{D_{0}=2\}|D_{0}\neq 1,D_{1}\neq 1\}=1~{},$ (16)
$Q\{\mathbbm{1}\{D_{2}=1\}$ $=\mathbbm{1}\{D_{0}=1\}|D_{0}\neq 2,D_{2}\neq 2\}=1~{}.$ (17)
The condition in (16) states that if crossing the cutoff for field 1 does not cause the student to switch to field 1, then it does not cause them to switch to or away from field 2. A similar interpretation applies to (17). Lee and Salanié (2023) show the set of all distributions that satisfy (14)–(17) are equivalent to a strict one-to-one targeting model with $|\mathcal{Z}|=3$ and $|\mathcal{D}^{\dagger}|=2$; it therefore follows from Remark 4.7 that any $Q$ that satisfies (14)–(17) also satisfies Assumption 2.2. Here, we establish directly that (14)–(17) imply Assumption 2.2 with $z^{\ast}(0)=0$, $z^{\ast}(1)=1$, and $z^{\ast}(2)=2$. To show $z^{\ast}(0)=0$, we prove by contradiction that
$Q\{D_{0}\neq 0,D_{1}=0\}=0~{}.$
Suppose with positive probability that $D_{0}\neq 0$ but $D_{1}=0$. On this event, (14) implies $D_{0}\neq 1$, so $D_{0}=2$. But $D_{1}=0$, which contradicts (16). Similarly,
$Q\{D_{0}\neq 0,D_{2}=0\}=0~{},$
and therefore $z^{\ast}(0)=0$. To show $z^{\ast}(1)=1$, first note (14) implies
$Q\{D_{1}\neq 1,D_{0}=1\}=0~{}.$
It therefore remains to argue by contradiction that
$Q\{D_{1}\neq 1,D_{2}=1\}=0~{}.$ (18)
Suppose with positive probability that $D_{1}\neq 1$ but $D_{2}=1$. On this event, (14) implies $D_{0}\neq 1$. If $D_{0}=2$, then (15) implies $D_{2}=2$, a contradiction to $D_{2}=1$; if instead $D_{0}=0$, then because we assume $D_{2}=1$, (17) implies $D_{2}\neq 1$, another contradiction. Therefore, (18) holds, and $z^{\ast}(1)=1$. $z^{\ast}(2)=2$ can be established following similar arguments.
5 Examples of Models That Do Not Satisfy Assumption 2.2
We now consider models that do not satisfy generalized monotonicity (Assumption 2.2). For each model, we show that the identified sets for average potential outcomes are not given by (5). For the first two examples, we further show that they do in fact provide identifying power beyond instrument exogeneity.
Example 5.1.
Suppose $\mathcal{Y}=\mathcal{D}=\mathcal{Z}=\{0,1\}$. Let $\mathbf{Q}$ denote all distributions $Q$ that satisfy Assumption 2.1 and
$Q\{D_{0}=D_{1}\}=0~{},$
which, in the language of Angrist et al. (1996), is imposing that all individuals are either compliers or defiers. For any $P$ such that $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$, $\theta(Q)$ is identified relative to $\mathbf{Q}$, i.e., $\Theta_{0}(P,\mathbf{Q})$ is a singleton. To see this, note that for any $Q\in\mathbf{Q}_{0}(P,\mathbf{Q})$ and $y\in\mathcal{Y}$,
$Q\{Y_{1}=y\}$ $=Q\{Y_{1}=y,D_{0}=0,D_{1}=1\}+Q\{Y_{1}=y,D_{0}=1,D_{1}=0\}$
$=Q\{Y_{1}=y,D_{1}=1\}+Q\{Y_{1}=y,D_{0}=1\}$
$=Q\{Y_{1}=y,D_{1}=1|Z=1\}+Q\{Y_{1}=y,D_{0}=1|Z=0\}$
$=P\{Y=y,D=1|Z=1\}+P\{Y=y,D=1|Z=0\}~{},$
where the first two equalities follows from $Q\{D_{0}=D_{1}\}=0$, the third equality follows from Assumption 2.1, and the final equality follows from $Q\in\mathbf{Q}_{0}(P,\mathbf{Q})$. A similar argument establishes identification of $Q\{Y_{0}=y\}$. In contrast, we show in Appendix B.2 that there exists a $P$ for which $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$ and (5) is not a singleton. The identified set for $\theta(Q)$ is therefore not given by (5). We further show in Appendix B.2 that $\Theta_{0}(P,\mathbf{Q}_{1}^{\ast})$ is not a singleton for the same $P$. Thus, the model $\mathbf{Q}$ does have identifying power beyond instrument exogeneity for $\theta(Q)$.
Example 5.2.
Consider an ordered choice model for treatment. Suppose that $|\mathcal{Z}|\geq 3$, and let $\mathbf{Q}$ denote the set of all distributions that satisfy Assumption 2.1 and
$Q\{D_{j}\geq D_{k}\}=1\text{ for all }j\geq k~{}.$ (19)
For example, $D$ might represent quantity of some treatment, and $Z$ might represent levels of subsidy for the treatment. The restriction in (19) is equivalent to the monotonicity assumption considered in Angrist and Imbens (1995). See Vytlacil (2006) for the connection between this restriction and ordered discrete-choice selection models. Without loss of generality, let $\mathcal{D}=\{0,\ldots,\bar{D}\}$. In this case, (3) is satisfied for $d\in\{0,\bar{D}\}$ for all $Q\in\mathbf{Q}_{0}(P,\mathbf{Q})$; to see this, take $z^{*}(0)=\min\{\mathcal{Z}\}$ and $z^{*}(\bar{D})=\max\{\mathcal{Z}\}$. By a straightforward modification of the arguments underlying Theorem 3.2, one can show that the identified sets for $\mathbb{E}_{Q}[Y_{0}]$ and $\mathbb{E}_{Q}[Y_{\bar{D}}]$ are given by (5) for any $P$ such that $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$. The ordered monotonicity assumption in (19) therefore has no identifying power beyond instrument exogeneity for $\mathbb{E}_{Q}[Y_{0}]$ and $\mathbb{E}_{Q}[Y_{\bar{D}}]$. In contrast, (3) need not hold for $d\in\mathcal{D}\setminus\{0,\bar{D}\}$ and $Q\in\mathbf{Q}_{0}(P,\mathbf{Q})$. In Appendix B.3, we show there exists a $P$ for which $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$ and $\Theta_{0}(P,\mathbf{Q})$ is not given by (5). We further show $\Theta_{0}(P,\mathbf{Q})\subsetneqq\Theta_{0}(P,\mathbf{Q}_{1}^{\ast})$ for the same $P$. Thus, the ordered monotonicity assumption in (19) does have identifying power beyond instrument exogeneity for $\mathbb{E}_{Q}[Y_{d}]$ for $d\in\mathcal{D}\setminus\{0,\bar{D}\}$.
Example 5.3.
In Example 4.7, we considered ARUMs satisfying the strict one-to-one targeting assumption of Lee and Salanié (2023). As discussed there, if $|\mathcal{D}|=2$ or if $|\mathcal{D}|\geq 3$ and $|\mathcal{Z}|>|\mathcal{D}^{\dagger}|$, then Assumption 2.2 holds and the results in Section 3 are applicable. Now consider the case in which $|\mathcal{D}|\geq 3$ and $|\mathcal{Z}|=|\mathcal{D}^{\dagger}|$. Denote by $\mathbf{Q}$ the ARUM model defined by (10) under the additional assumption that the support of $(U_{d}:d\in\mathcal{D})$ is $\Re^{|\mathcal{D}|}$. Then, as we show in Appendix B.4, while (3) will hold for targeted treatments, (3) cannot hold for any non-targeted treatment, and thus Assumption 2.2 is violated. We show that, while the identified set for $\mathbb{E}_{Q}[Y_{d}]$ is given by (5) for targeted treatments when $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$, there exists a $P$ for which $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$ and the identified set for $\mathbb{E}_{Q}[Y_{d}]$ is not given by (5) for non-targeted treatments.
Appendix A Proofs of Main Results
A.1 Proof of Lemma 2.1
The necessity of (4) has been proved in the main text. On the other hand, fix a $d\in\mathcal{D}$, suppose (4) holds for some $z\in\mathcal{Z}$. Fix a particular $z^{\ast}\in\mathcal{Z}^{*}(d,Q)$ that satisfies (3). If $z=z^{\ast}$ then of course Assumption 2.2 holds for $z$. Suppose $z\neq z^{\ast}$. Then,
$0\leq$ $~{}P\{D=d\mid Z=z\}-P\{D=d\mid Z=z^{\ast}\}$
$=$ $~{}Q\{D_{z}=d\}-Q\{D_{z^{\ast}}=d\}$
$=$ $~{}Q\{D_{z^{\ast}}\neq d,D_{z}=d\}-Q\{D_{z}\neq d,D_{z^{\ast}}=d\}$
$=$ $~{}-Q\{D_{z}\neq d,D_{z^{\ast}}=d\}~{},$
where the first inequality is using that $z$ satisfies (4), the second line is using Assumption 2.1, and the last line is using that $z^{\ast}$ satisfies (3). Thus $Q\{D_{z}\neq d,D_{z^{\ast}}=d\}=0$. Assumption 2.2 implies
$Q\{D_{z}\neq d,D_{z^{\prime}}=d\text{ for some }z^{\prime}\in\mathcal{Z}\}$
$=Q\{D_{z}\neq d,D_{z^{\ast}}=d\}+Q\{D_{z}\neq d,D_{z^{\ast}}\neq d,D_{z^{\prime}}=d~{}\text{for some}~{}z^{\prime}\in\mathcal{Z}\setminus\{z^{\ast}\}\}$
$=0~{},$
and $z$ satisfies Assumption 2.2 as well.
A.2 Proof of Theorem 3.1
The desired result follows immediately from Theorems 3.2 and 3.3.
A.3 Proof of Theorem 3.2
A.3.1 Auxillary Results
To present the proof of Theorem 3.2, we first introduce some further notation. Recall from the discussion in Section 2 that if $Q$ satisfies Assumption 2.1, then $P=QT^{-1}$ if and only if
$p_{yd|z}=Q\{Y_{d}=y,D_{z}=d\}~{}.$
Following Heckman and Pinto (2018), we define a treatment response type as a vector $r^{t}\in\mathcal{D}^{|\mathcal{Z}|}$,
$r^{t}$ $=(d_{z}:z\in\mathcal{Z})\in\mathcal{D}^{|\mathcal{Z}|}~{}.$
Treatment response types are also called principal strata (Frangakis and Rubin, 2002). We analogously define an outcome response type as a vector $r^{o}\in\mathcal{Y}^{|\mathcal{D}|}$,
$r^{o}$ $=(y_{d}:d\in\mathcal{D})\in\mathcal{Y}^{|\mathcal{D}|}~{}.$
Let $r_{j}^{o}$ denote the $(j+1)$th entry of $r^{o}$ and $r_{j}^{t}$ denote the $(j+1)$th entry of $r^{t}$. In this notation,
$Q\{Y_{d}=y,D_{z}=d\}=\sum_{(r^{o},r^{t}):r^{o}_{d}=y,r^{t}_{z}=d}q(r^{o},r^{t})~{},$
where
$q\big{(}r^{o},r^{t}\big{)}=Q\{(Y_{d}:d\in\mathcal{D})=r^{o},(D_{z}:z\in\mathcal{Z})=r^{t}\}~{}.$
Thus, if $Q$ satisfies Assumption 2.1, then $P=QT^{-1}$ if and only if
$p_{yd|z}=\sum_{(r^{o},r^{t}):r^{o}_{d}=y,r^{t}_{z}=d}q(r^{o},r^{t})\qquad\forall~{}y\in\mathcal{Y},~{}d\in\mathcal{D},~{}z\in\mathcal{Z}~{}.$ (20)
We note Assumption 2.2 can be expressed as restricting $q(r^{o},r^{t})=0$ for certain $r^{t}$ and $r^{o}$. In particular, if for some $d\in\mathcal{D}$, $r^{t}_{z^{*}(d)}\neq d$ while $r^{t}_{z^{\prime}}=d$ for some $z^{\prime}\neq z^{*}(d)$, then $q(r^{o},r^{t})=0$ for all $r^{o}$. Let $R^{t}$ denote the complement of that set of treatment response types, or equivalently the set of treatment response types that may occur with positive probability under Assumption 2.2:
$R^{t}=\{r^{t}\in\mathcal{D}^{|\mathcal{Z}|}~{}:~{}r^{t}_{z^{*}(d)}\neq d~{}\mbox{implies}~{}r^{t}_{z^{\prime}}\neq d~{}\forall~{}d\in\mathcal{D},z^{\prime}\in\mathcal{Z}\}~{}.$ (21)
If Assumption 2.2 is satisfied, then $Q\{(D_{z}:z\in\mathcal{Z})\in R^{t}\}=1$.
Below we derive a lemma that simplifies determining whether $q(r^{o},r^{t})$ satisfies (20) and will be used subsequently to derive our characterization of the identified set. To this end, we require some further notation. Let
$\mathcal{N}(r^{t})$ =$=\{d\in\mathcal{D}~{}:~{}r^{t}_{z}\neq d~{}\mbox{for all}~{}z\in\mathcal{Z}\}~{},$
$\mathcal{N}(r^{t})^{c}$ =$=\{d\in\mathcal{D}~{}:~{}r^{t}_{z}=d~{}\mbox{for some}~{}z\in\mathcal{Z}\}~{},$
For a given treatment response type $r^{t}$, $\mathcal{N}(r^{t})$ is the set of treatments for which that treatment response type is a never-taker, and $\mathcal{N}(r^{t})^{c}$ is the set of treatments for which that treatment response type will take that treatment for some value of $z$. Using this notation, partition outcome and treatment response types $(r^{o},r^{t})$ as $(r^{o}_{n}(r^{t}),r^{o}_{c}(r^{t}),r^{t})$ where
$r^{o}_{n}(r^{t})$ $=(r^{o}_{d}:d\in\mathcal{N}(r^{t}))~{},$
$r^{o}_{c}(r^{t})$ $=(r^{o}_{d}:d\in\mathcal{N}(r^{t})^{c})~{}.$
For a given treatment response type $r^{t}$, $r^{o}_{n}(r^{t})$ are those outcomes that are never observed for that response type, and $r^{o}_{c}(r^{t})$ are the remaining outcomes that are observed given some potential value of $Z$. For example, for any $r^{t}\in R^{t}$ with $R^{t}$ defined by (21),
$\mathcal{N}(r^{t})$ =$=\{d\in\mathcal{D}~{}:~{}r^{t}_{z^{*}(d)}\neq d\}~{},$
$\mathcal{N}(r^{t})^{c}$ =$=\{d\in\mathcal{D}~{}:~{}r^{t}_{z^{*}(d)}=d\}~{},$
and
$r^{o}_{n}(r^{t})$ $=(r^{o}_{d}:d\in\mathcal{D},~{}r^{t}_{z^{*}(d)}\neq d)~{},$
$r^{o}_{c}(r^{t})$ $=(r^{o}_{d}:d\in\mathcal{D},~{}r^{t}_{z^{*}(d)}=d)~{}.$
For notational convenience, further define $q(r^{o}_{c}(r^{t}),r^{t})$ as $q(r^{o}_{n}(r^{t}),r^{o}_{c}(r^{t}),r^{t})$ summed over $r^{o}_{n}(r^{t})$:
$q(r^{o}_{c}(r^{t}),r^{t})=\begin{cases}q(r^{o},r^{t})&\mbox{if}~{}\mathcal{N}(r^{t})=\emptyset~{}\mbox{so that}~{}r^{o}_{c}(r^{t})=r^{o}\\ \sum_{r^{o}_{n}(r^{t})\in\mathcal{Y}^{|\mathcal{N}(r^{t})|}}q(r^{o}_{n}(r^{t}),r^{o}_{c}(r^{t}),r^{t})&\mbox{if}~{}\mathcal{N}(r^{t})\neq\emptyset~{}\mbox{so that}~{}r^{o}_{c}(r^{t})\neq r^{o}~{}.\end{cases}$
Using this notation, we have the following lemma that asserts whether $q(r^{o},r^{t})$ satisfies (20) depends only on $q(r^{o}_{c}(r^{t}),r^{t})$.
Lemma A.1.
Suppose $q$ satisfies (20). Then, $q^{*}$ satisfies (20) if, for each $r^{t}\in\mathcal{D}^{|\mathcal{Z}|}$,
$q^{*}(r^{o}_{c}(r^{t}),r^{t})=q(r^{o}_{c}(r^{t}),r^{t})~{}~{}\forall~{}r^{o}_{c}(r^{t})~{}.$
Proof.
We can rewrite (20) as
$p_{yd|z}=$ $~{}\sum_{r^{t}:r^{t}_{z}=d}~{}~{}\sum_{r^{o}:r^{o}_{d}=y}q(r^{o},r^{t})$
$=$ $\sum_{r^{t}:r^{t}_{z}=d}~{}~{}\sum_{r^{o}_{c}(r^{t}):r^{o}_{d}=y}~{}~{}\left(\sum_{r^{o}_{n}(r^{t})}q(r^{o}_{n}(r^{t}),r^{o}_{c}(r_{t}),r^{t})\right)$
$=$ $\sum_{r^{t}:r^{t}_{z}=d}~{}~{}\sum_{r^{o}_{c}(r^{t}):r^{o}_{d}=y}~{}~{}q(r^{o}_{c}(r_{t}),r^{t})~{},$
where the second equality uses that $r_{c}^{o}(r^{t})$ is nonempty because $r^{t}_{z}=d$ and that $r^{o}_{d}$ is an element of $r^{o}_{c}(r^{t})$ for $r^{t}$ such that $r^{t}_{z}=d$. The result now follows.
A.3.2 Proof of the Theorem
Suppose $Q\in\mathbf{Q}_{0}(P,\mathbf{Q})$. For each $d\in\mathcal{D}$,
$\mathbb{E}_{Q}[Y_{d}]$ $=\mathbb{E}_{Q}[Y_{d}\mathbbm{1}\{D_{z^{*}(d)}=d\}]+\mathbb{E}_{Q}[Y_{d}\mathbbm{1}\{D_{z^{*}(d)}\neq d\}]$
$=\beta_{d|z^{*}(d)}+\mathbb{E}_{Q}[Y_{d}\mathbbm{1}\{D_{z^{*}(d)}\neq d\}]~{},$
where the second equality is using Assumption 2.1. We have
$\mathbb{E}[Y_{d}\mathbbm{1}\{D_{z^{*}(d)}\neq d\}]\in[y^{L}Q\{D_{z^{*}(d)}\neq d\},~{}y^{U}Q\{D_{z^{*}(d)}\neq d\}]~{},$
while Assumption 2.1 implies that
$Q\{D_{z^{*}(d)}\neq d\}=1-\sum_{y\in\mathcal{Y}}p_{yd|z^{*}(d)}~{}.$
We thus have that
$\mathbb{E}[Y_{d}]\in[\beta_{d|z^{*}(d)}+y^{L}(1-\sum_{y\in\mathcal{Y}}p_{yd|z^{*}(d)}),~{}\beta_{d|z^{*}(d)}+y^{U}(1-\sum_{y\in\mathcal{Y}}p_{yd|z^{*}(d)})]~{},$
for each $d\in\mathcal{D}$, and thus (5) provides valid bounds on $\theta(Q)$ under the stated assumptions.
Let $q$ denote latent probabilities corresponding to a fixed $Q\in\mathbf{Q}_{0}(P,\mathbf{Q})$, which is nonempty by assumption. We now show that for any value in (5) there exists $Q^{\ast}\in\mathbf{Q}_{0}(P,\mathbf{Q})$ for which $(\mathbb{E}_{Q^{\ast}}[Y_{j}]:j\in\mathcal{D})$ equals the prescribed value. To do so, we show that, for each $\theta$ in the right-hand side of (5), we can construct $q^{\ast}$ that 1) is nonnegative and has $\sum_{r^{o}}q^{\ast}(r^{o},r^{t})=\sum_{r^{o}}q(r^{o},r^{t})$, 2) satisfies (20) and hence $Q^{\ast}\in\mathbf{Q}_{0}(P,\mathbf{Q})$, and 3) satisfies $\theta({Q^{\ast}})=\theta$.
Fix some $\alpha=(\alpha_{0},\alpha_{1},...,\alpha_{|\mathcal{D}|-1})^{\prime}\in[0,1]^{|\mathcal{D}|}$ to be specified below, we now construct an alternative $q^{\ast}$ as follows. For each treatment response type $r^{t}$, let $q^{\ast}_{0}(r^{o},r^{t})=q(r^{o},r^{t})$ for all $r^{o}$. Let $K(r^{t})=|\mathcal{N}(r^{t})|$. Note that if $K(r^{t})=0$, then $\mathcal{N}(r^{t})=\emptyset$, so $r^{t}_{z^{*}(d)}=d$ for all $d\in\mathcal{D}$ under $r^{t}$. For such an $r^{t}$, we set $q^{\ast}(r^{o},r^{t})=q(r^{o},r^{t})$ for all $r^{o}\in\mathcal{Y}^{|\mathcal{D}|}$. If $K(r^{t})\geq 1$, enumerate $\mathcal{N}(r^{t})$ as $\{j[1],...,j[K(r^{t})]\}$, and for $k=1$ to $K(r^{t})$, define $q_{k}^{\ast}$ iteratively as follows:
$\begin{split}q^{*}_{k}((r^{o}_{-j[k]},r^{o}_{j[k]}=y^{L}),r^{t})&=(1-\alpha_{j[k]})\sum_{r^{o}_{j[k]}\in\mathcal{Y}}q^{*}_{k-1}((r^{o}_{-j[k]},r^{o}_{j[k]}),r^{t})\\ q^{*}_{k}((r^{o}_{-j[k]},r^{o}_{j[k]}=y),r^{t})&=0\qquad\mbox{for}~{}y\not\in\{y^{L},y^{U}\}\\ q^{\ast}_{k}((r^{o}_{-j[k]},r^{o}_{j[k]}=y^{U}),r^{t})&=\alpha_{j[k]}\sum_{r^{o}_{j[k]}\in\mathcal{Y}}q^{*}_{k-1}((r^{o}_{-j[k]},r^{o}_{j[k]}),r^{t})~{},\end{split}$ (22)
for all $r^{o}_{-j[k]}$, where we partition $r^{o}=(r^{o}_{-j[k]},r^{o}_{j[k]})$. With this construction, the marginal distribution of $Y_{j[k]}$ for treatment response type $r^{t}$ is only modified in step $k$. In particular, for each fixed $k$, for $\ell\leq k-1$, for $y\in\mathcal{Y}$,
$\sum_{r_{-j[k]}^{o}}q^{*}_{\ell}((r^{o}_{-j[k]},r^{o}_{j[k]}=y),r^{t})=\sum_{r_{-j[k]}^{o}}q^{*}_{0}((r^{o}_{-j[k]},r^{o}_{j[k]}=y),r^{t})=\sum_{r_{-j[k]}^{o}}q((r^{o}_{-j[k]},r^{o}_{j[k]}=y),r^{t})~{},$ (23)
which together with (22) implies
$\mathbb{E}_{Q^{\ast}_{k}}[Y_{j[k]}\mathbbm{1}\{(D_{z}:z\in\mathcal{Z})=r^{t}\}]$ $=\sum_{y\in\mathcal{Y}}\sum_{r_{-j[k]}^{o}}yq^{*}_{k}((r^{o}_{-j[k]},r^{o}_{j[k]}=y),r^{t})$
$=(\alpha_{j[k]}y^{U}+(1-\alpha_{j[k]})y^{L})\sum_{r^{o}}q(r^{o},r^{t})~{}.$
On the other hand, for each fixed $k$, because the marginal distribution of $Y_{j[k]}$ for treatment response type $r^{t}$ is not further modified after step $k$, (22) and (23) imply
$\begin{split}\sum_{r_{-j[k]}^{o}}q^{*}_{K(r^{t})}((r^{o}_{-j[k]},r^{o}_{j[k]}=y^{L}),r^{t})&=\sum_{r_{-j[k]}^{o}}q^{*}_{k}((r^{o}_{-j[k]},r^{o}_{j[k]}=y^{L}),r^{t})\\ &=(1-\alpha_{j[k]})\sum_{r^{o}_{j[k]}\in\mathcal{Y}}\sum_{r_{-j[k]}^{o}}q((r^{o}_{-j[k]},r^{o}_{j[k]}),r^{t})\\ &=(1-\alpha_{j[k]})\sum_{r^{o}}q(r^{o},r^{t})~{},\\ \sum_{r_{-j[k]}^{o}}q^{*}_{K(r^{t})}((r^{o}_{-j[k]},r^{o}_{j[k]}=y),r^{t})&=0\qquad\mbox{for}~{}y\not\in\{y^{L},y^{U}\}~{},\\ \sum_{r_{-j[k]}^{o}}q^{*}_{K(r^{t})}((r^{o}_{-j[k]},r^{o}_{j[k]}=y^{U}),r^{t})&=\sum_{r_{-j[k]}^{o}}q^{*}_{k}((r^{o}_{-j[k]},r^{o}_{j[k]}=y^{U}),r^{t})\\ &=\alpha_{j[k]}\sum_{r^{o}_{j[k]}\in\mathcal{Y}}\sum_{r_{-j[k]}^{o}}q((r^{o}_{-j[k]},r^{o}_{j[k]}),r^{t})\\ &=\alpha_{j[k]}\sum_{r^{o}}q(r^{o},r^{t})~{}.\end{split}$ (24)
Set
$q^{*}(r^{o},r^{t})=q^{*}_{K(r^{t})}(r^{o},r^{t})\qquad\forall~{}r^{o}~{}.$
With this construction, $q^{*}(r^{o},r^{t})$ is non-negative and
$\sum_{r^{o}}q^{*}(r^{o},r^{t})=\sum_{r^{o}}q(r^{o},r^{t})~{}~{}\mbox{for all}~{}~{}r^{t}~{}.$
Note that, since Assumption 2.2 holds for $Q$, $r^{o}_{d}$ is a component of $r_{c}^{o}(r^{t})$ for $r^{t}$ such that $r^{t}_{z^{*}(d)}=d$, while $r^{o}_{d}$ is a component of $r_{n}^{o}(r^{t})$ for $r^{t}$ such that $r^{t}_{z^{*}(d)}\neq d$. Therefore, with this construction, for each $r^{t}$, $q^{*}(r^{o}_{c}(r^{t}),r^{t})=q(r^{o}_{c}(r^{t}),r^{t})~{}~{}\forall~{}r^{o}_{c}(r^{t})$ because of (24), and hence, by Lemma A.1, $q^{*}$ satisfies (20) so that $Q^{\ast}\in\mathbf{Q}_{0}(P,\mathbf{Q})$. That $q^{*}$ satisfies (20) implies that, for each $d\in\mathcal{D}$,
$\mathbb{E}_{Q^{*}}[Y_{d}\mathbbm{1}\{D_{z^{*}(d)}=d\}]=\beta_{d|z^{*}(d)}~{}.$
We also have that for each $d\in\mathcal{D}$,
$\mathbb{E}_{Q^{\ast}}[Y_{d}\mathbbm{1}\{D_{z^{*}(d)}\neq d\}]$ $=\sum_{y\in\mathcal{Y}}~{}~{}\sum_{r^{t}:r^{t}_{z^{*}(d)}\neq d}~{}~{}\sum_{r^{o}:r^{o}_{d}=y}~{}~{}y~{}q^{*}(r^{o},r^{t})$
$=\sum_{y\in\mathcal{Y}}~{}~{}\sum_{r^{t}:r^{t}_{z^{*}(d)}\neq d}~{}~{}\sum_{r^{o}_{-d}}~{}~{}y~{}q^{*}((r^{o}_{-d},r^{o}_{d}=y),r^{t})$
$=(\alpha_{d}y^{U}+(1-\alpha_{d})y^{L})\sum_{r^{t}:r^{t}_{z^{*}(d)}\neq d}~{}~{}\sum_{r^{o}}q(r^{o},r^{t})$
$=(\alpha_{d}y^{U}+(1-\alpha_{d})y^{L})~{}Q\{D_{z^{*}(d)}\neq d\}$
$=(\alpha_{d}y^{U}+(1-\alpha_{d})y^{L})(1-\sum_{y}p_{yd|z^{*}(d)})~{},$
where the third equality is using that (24) holds for $r^{t}$ such that $r^{t}_{z^{*}(d)}\neq d$, so that $d=j[k^{\prime}]$ for some $k^{\prime}$ in constructing $q^{*}(\cdot,r^{t})$ for that $r^{t}$, and the last equality is using that $Q$ satisfies (20). Thus, for each $d\in\mathcal{D}$,
$\mathbb{E}_{Q^{*}}[Y_{d}]$ $=\mathbb{E}_{Q^{*}}[Y_{d}\mathbbm{1}\{D_{z^{*}(d)}=d\}]+\mathbb{E}_{Q^{*}}[Y_{d}\mathbbm{1}\{D_{z^{*}(d)}\neq d\}]$
$=\beta_{d|z^{*}(d)}+(\alpha_{d}y^{U}+(1-\alpha_{d})y^{L})(1-\sum_{y}p_{yd|z^{*}(d)})~{}.$
For any $\theta$ contained in (5), we can thus chose $\alpha=(\alpha_{0},\alpha_{1},...,\alpha_{|\mathcal{D}|-1})^{\prime}\in[0,1]^{|\mathcal{D}|}$ such that $\theta({Q^{\ast}})=\theta$.
A.4 Proof of Corollary 3.1
The results follows immediately from Theorem 3.2 because $\mathbb{E}[Y_{j}]-\mathbb{E}[Y_{k}]$ is a function of $\theta(Q)$.
A.5 Proof of Lemma 3.1
To see it, note for any $d\in\mathcal{D}$ and $z\in\mathcal{Z}$,
$\mathbb{E}_{Q}[Y_{d}]=\mathbb{E}_{Q}[Y_{d}|Z=z]$ $=\mathbb{E}_{Q}[Y_{d}\mathbbm{1}\{D=d\}|Z=z]+\mathbb{E}_{Q}[Y_{d}\mathbbm{1}\{D\neq d\}|Z=z]$
$=\mathbb{E}_{Q}[Y\mathbbm{1}\{D=d\}|Z=z]+\mathbb{E}_{Q}[Y_{d}\mathbbm{1}\{D\neq d\}|Z=z]$
$=\beta_{d|z}+\mathbb{E}_{Q}[Y_{d}\mathbbm{1}\{D\neq d\}|Z=z]$
$\leq\beta_{d|z}+y^{U}P\{D\neq d|Z=z\}$
$=\beta_{d|z}+y^{U}(1-\sum_{y\in\mathcal{Y}}p_{yd|z})~{}.$
Because the inequality holds for all $z\in\mathcal{Z}$, the upper end for each $d\in\mathcal{D}$ of (7) is a valid upper bound for $\mathbb{E}[Y_{d}]$. On the other hand, they can be simultaneously attained for all $d\in\mathcal{D}$ by setting $Y_{d}=y^{U}$ whenever $D\neq d$ and $Z=z$, without affecting the distribution of $(Y,D,Z)$. A similar argument can be applied to the lower ends. In addition, any values in between can also be attained simultaneously for all $d\in\mathcal{D}$ by setting $Y_{d}$ to be a convex combination of $y^{L}$ and $y^{U}$ whenever $D\neq d$ and $Z=z$ without affecting the distribution of $(Y,D,Z)$, and therefore (7) is indeed the identified set for $\theta(Q)$ under mean independence.
A.6 Proof of Lemma 3.2
Suppose $Q\in\mathbf{Q}_{0}(P,\mathbf{Q}_{1,2}^{*})\neq\emptyset$. Consider the upper endpoints of (5) and (7). For any $d\in\mathcal{D}$, $z\in\mathcal{Z}$,
$\mathbb{E}_{P}[Y$ $\mathbbm{1}\{D=d\}]\mid Z=z^{*}(d)]+y^{U}\mathbb{E}_{P}[1-\mathbbm{1}\{D=d\}\mid Z=z^{*}(d)]$
$~{}~{}~{}~{}-\mathbb{E}_{P}[Y\mathbbm{1}\{D=d\}\mid Z=z]-y^{U}\mathbb{E}_{P}[1-\mathbbm{1}\{D=d\}\mid Z=z]$
$=\mathbb{E}_{Q}[Y_{d}\mathbbm{1}\{D_{z^{*}(d)}=d\}]+y^{U}\mathbb{E}_{Q}[1-\mathbbm{1}\{D_{z^{*}(d)}=d\}]-\mathbb{E}_{Q}[Y_{d}\mathbbm{1}\{D_{z}=d\}]-y^{U}\mathbb{E}_{Q}[1-\mathbbm{1}\{D_{z}=d\}]$
$=\mathbb{E}_{Q}[(Y_{d}-y^{U})(\mathbbm{1}\{D_{z^{*}(d)}=d\}-\mathbbm{1}\{D_{z}=d\})]$
$=\mathbb{E}_{Q}[(Y_{d}-y^{U})(\mathbbm{1}\{D_{z^{*}(d)}=d,D_{z}\neq d\}-\mathbbm{1}\{D_{z^{*}(d)}\neq d,D_{z}=d\})]$
$=\mathbb{E}_{Q}[(Y_{d}-y^{U})\mathbbm{1}\{D_{z^{*}(d)}=d,D_{z}\neq d\}]$
$\leq 0~{},$
where the first equality uses Assumption 2.1 and the fourth equality uses that $Q\{D_{z^{*}(d)}\neq d,D_{z}=d\}=0$ for all $Q$ satisfying Assumption 2.2. Since this inequality holds for all $z\in\mathcal{Z}$, we have that the upper endpoint of the interval in (5) is weakly smaller than the upper endpoint of the interval in (7). Conversely, the upper endpoint of (5) is contained in the set over which the upper endpoint of (7) is minimizing over, and thus the upper endpoint of (5) is weakly larger than the upper endpoint of (7). We conclude that the upper endpoints are the same. Parallel arguments show equivalence of the lower endpoints.
A.7 Proof of Theorem 3.3
Recall that Assumption 2.1 holds for $Q\in\mathbf{Q}_{0}(P,\mathbf{Q}_{1,2}^{*})$. Note that Assumption 2.1 implies Assumption 3.1, and hence $\mathbf{Q}_{1,2}^{\ast}\subseteq\mathbf{Q}_{1}^{\ast}\subseteq\mathbf{Q}_{3}^{\ast}$. As a result,
$\Theta_{0}(P,\mathbf{Q}_{1,2}^{*})\subseteq\Theta_{0}(P,\mathbf{Q}_{1}^{*})\subseteq\Theta_{0}(P,\mathbf{Q}_{3}^{*})~{}.$
By Lemma 3.2, when $\mathbf{Q}_{0}(P,\mathbf{Q}_{1,2}^{*})\neq\emptyset$, we have
$\Theta_{0}(P,\mathbf{Q}_{1,2}^{*})=\Theta_{0}(P,\mathbf{Q}_{3}^{*})~{}.$
The result now follows by a sandwich argument.
Appendix B Details of Examples
B.1 Details of Example 4.7
We show that, except in the special case where the treatment and the instrument are both binary, the strict one-to-one targeting assumption of Lee and Salanié (2023) with one or more targeted treatments implies that (11) does not hold for some treatments. To see this, suppose that the strict one-to-one targeting assumption of Lee and Salanié (2023) holds with $|\mathcal{D}^{\dagger}|\geq 1$. For each $d\in\mathcal{D}^{\dagger}$, their assumptions include that there exists some $z^{\dagger}(d)$ and some $\overline{U}(d)$, $\underline{U}(d)$ with $\overline{U}(d)>\underline{U}(d)$ such that
$g(z,d)=$$\begin{cases}\overline{U}(d)&\mbox{if}~{}z=z^{\dagger}(d)\\ \underline{U}(d)&\mbox{if}~{}z\neq z^{\dagger}(d)~{}.\end{cases}$ (25)
On the other hand, for each $d\in\mathcal{D}\setminus\mathcal{D}^{\dagger}$, they impose that $g(z,d)=\underline{U}(d)$ for all $z\in\mathcal{Z}$, and they impose that there is at least one such non-targeted treatment. For any $d\in\mathcal{D}\setminus\mathcal{D}^{\dagger}$, (11) requires that there exists $z^{\ast}(d)\in\mathcal{Z}$ such that
$g(z,d^{\prime})>g(z^{\ast}(d),d^{\prime})~{}~{}\text{ for all }d^{\prime}\neq d\text{ and }z\neq z^{\ast}(d)~{}.$ (26)
Suppose $|\mathcal{Z}|\geq 3$, and fix some targeted treatment $d^{\prime}\in\mathcal{D}^{\dagger}$. Suppose $z^{\ast}(d)\neq z^{\dagger}(d^{\prime})$. Then, for $z\in\mathcal{Z}\setminus\{z^{\ast}(d),z^{\dagger}(d^{\prime})\}$, (26) requires $\underline{U}(d^{\prime})>\underline{U}(d^{\prime})$, a contradiction. Now suppose $z^{\ast}(d)=z^{\dagger}(d^{\prime})$. Then, for $z\in\mathcal{Z}\setminus\{z^{\ast}(d)\}$, (26) requires $\underline{U}(d^{\prime})>\overline{U}(d^{\prime})$, a contradiction. Thus, $|\mathcal{Z}|\geq 3$ implies that (11) does not hold for non-targeted treatments.
Now suppose $|\mathcal{Z}|=2$, which we label as $\mathcal{Z}=\{0,1\}$, and suppose $|\mathcal{D}|\geq 3$. Without loss of generality suppose $1\in\mathcal{D}^{\dagger}$ and $z^{\dagger}(1)=1$. If $z^{\ast}(d)=1$, then (26) requires $\underline{U}(1)>\overline{U}(1)$, a contradiction. Now suppose $z^{\ast}(d)=0$. Then (26) requires $g(1,d^{\prime})>g(0,d^{\prime})$ for all $d^{\prime}\neq d$. Consider the following two cases:
  • If $|\mathcal{D}^{\dagger}|=1$, then $g(1,d^{\prime})>g(0,d^{\prime})$ holding for any $d^{\prime}\in(\mathcal{D}\setminus\mathcal{D}^{\dagger})\setminus\{d\}$ requires $\underline{U}(d^{\prime})>\underline{U}(d^{\prime})$, a contradiction.
  • If $|\mathcal{D}^{\dagger}|>1$, then there exists $d^{\prime\prime}\in D^{\dagger}\setminus\{1\}$. By assumption $z^{\dagger}(d^{\prime\prime})\neq z^{\dagger}(1)$ so that $z^{\dagger}(d^{\prime\prime})=0$. Then $g(1,d^{\prime})>g(0,d^{\prime})$ holding for $d^{\prime}=d^{\prime\prime}$ requires $\underline{U}(d^{\prime\prime})>\overline{U}(d^{\prime\prime})$, again a contradiction.
Thus, $|\mathcal{Z}|=2$ with $|\mathcal{D}|\geq 3$ implies that (11) does not hold for some treatments.
We have shown (11) does not hold for some treatments when either $Z$ or $D$ takes at least three values. Now suppose $|\mathcal{D}|=|\mathcal{Z}|=2$. Let $D=0$ denote the nontargeted treatment and $D=1$ the targeted treatment, and let $z^{\dagger}(1)=1$. Consider $z^{\ast}(0)=0$ and $z^{\ast}(1)=1$. Then evaluating (11) at either $d=0$ or $d=1$ results in $\overline{U}(1)>\underline{U}(1),$ and thus (11) holds when $|\mathcal{D}|=|\mathcal{Z}|=2$. We conclude that the strict one-to-one targeting of Lee and Salanié (2023) implies that (11) does not hold for some $d\in\mathcal{D}$ except in the special case where $|\mathcal{D}|=|\mathcal{Z}|=2$.
We now show that the strict one-to-one targeting of Lee and Salanié (2023) implies that Assumption 2.2 holds when $|\mathcal{Z}|>|\mathcal{D}^{\dagger}|$. Let $\mathcal{Z}^{\dagger}\subseteq\mathcal{Z}$ denote the set of instruments that target some treatment,
$\mathcal{Z}^{\dagger}=\{z\in\mathcal{Z}:z=z^{\dagger}(d)~{}\mbox{for some}~{}d\in\mathcal{D}^{\dagger}\}~{}.$
Their strict one-to-one targeting assumption combined with $|\mathcal{Z}|>|\mathcal{D}^{\dagger}|$ implies that there are values of the instrument that do not target any treatment; in other words, $\mathcal{Z}^{\dagger}\subsetneqq\mathcal{Z}$. Following Lee and Salanié (2023), we label the treatment that is known not to be targeted as treatment 00, so that $g(z,0)=\underline{U}(0)$ for all $z\in\mathcal{Z}$, and impose their normalization that $\underline{U}(0)=0$. Consider (3) for $d=0$. Note that
$Q\{D_{z^{*}(0)}\neq 0,~{}D_{z^{\prime}}=0~{}\mbox{for some}~{}z^{\prime}\neq z^{*}(0)\}=~{}Q\Big{\{}\bigcup_{d^{*}\neq 0,z^{\prime}\neq z^{*}(0)}D_{z^{*}(0)}=d^{*},~{}D_{z^{\prime}}=0\Big{\}}~{}.$
We wish to investigate whether there exists some $z^{*}(0)\in\mathcal{Z}$ such that the above probability is zero. Consider $z^{*}(0)$ equal to any value in $\mathcal{Z}\setminus\mathcal{Z}^{\dagger}$, i.e., any value of the instrument that does not target any treatment. For any fixed $d^{*}\neq 0,z^{\prime}\neq z^{*}(0)$, consider the event $\{D_{z^{*}(0)}=d^{*},~{}D_{z^{\prime}}=0\}$. Since $z^{*}(0)$ does not target any treatment and thus does not target $d^{*}$, $D_{z^{*}(0)}=d^{*}$ implies
$U_{0}-U_{d^{*}}\leq\underline{U}(d^{*})~{}.$ (27)
If $z^{\prime}$ targets $d^{*}$, then $D_{z^{\prime}}=0$ implies
$U_{0}-U_{d^{*}}\geq\overline{U}(d^{*})~{}.$ (28)
If $z^{\prime}$ does not target $d^{*}$, then $D_{z^{\prime}}=0$ implies
$U_{0}-U_{d^{*}}\geq\underline{U}(d^{*})~{}.$ (29)
Thus, the event $\{D_{z^{*}(0)}=d^{*},~{}D_{z^{\prime}}=0\}$ either requires (27) and (28) to jointly hold, which is a contradiction since $\overline{U}(d^{*})>\underline{U}(d^{*})$ , or requires (27) and (29) to jointly hold, which is a zero probability event given our assumption that the distribution of $(U_{d}:d\in\mathcal{D})$ is absolutely continuous w.r.t. Lebesgue measure. Thus $Q\{D_{z^{*}(0)}\neq 0,~{}D_{z^{\prime}}=0~{}\mbox{for some}~{}z^{\prime}\neq z^{*}(0)\}$ is a probability of a finite union of zero probability events, and thus, by Boole’s inequality, equals zero so that (3) holds for $d=0$. A parallel argument shows that (3) holds for any non-targeted treatment, and related argument shows that (3) holds for any targeted treatment. Thus, under the strict one-to-one targeting of Lee and Salanié (2023), when there are more values of the instrument than targeted treatments, Assumption 2.2 holds even though (11) is violated for some $d\in\mathcal{D}$.
B.2 Details of Example 5.1
Let $\mathbf{Q}$ denote all distributions $Q$ for which Assumption 2.1 holds and such that $Q\{D_{0}=D_{1}\}=0$. Then for $Q\in\mathbf{Q}_{0}(P,\mathbf{Q})$,
$p_{y1|1}$ $=Q\{Y_{1}=y,D_{1}=1,D_{0}=0\}$
$p_{y0|0}$ $=Q\{Y_{0}=y,D_{1}=1,D_{0}=0\}$
$p_{y0|1}$ $=Q\{Y_{0}=y,D_{1}=0,D_{0}=1\}$
$p_{y1|0}$ $=Q\{Y_{1}=y,D_{1}=0,D_{0}=1\}$
and
$Q\{Y_{0}=1\}$ $=Q\{Y_{0}=1,D_{1}=1,D_{0}=0\}+Q\{Y_{0}=1,D_{1}=0,D_{0}=1\}$
$=p_{10|0}+p_{10|1}$
$Q\{Y_{1}=1\}$ $=Q\{Y_{1}=1,D_{1}=1,D_{0}=0\}+Q\{Y_{1}=1,D_{1}=0,D_{0}=1\}$
$=p_{11|1}+p_{11|0}~{}.$
Therefore, if $\mathbf{Q}$ is consistent with $P$, then $\theta(Q)$ is identified as
$\Theta_{0}(P,\mathbf{Q})=\left\{\begin{pmatrix}p_{10|0}+p_{10|1}\\ p_{11|0}+p_{11|1}\end{pmatrix}\right\}~{}.$ (30)
In contrast, the identified set that follows from imposing Assumption 2.1 alone, $\Theta_{0}(P,\mathbf{Q}_{1}^{\ast})$, is shown by Balke and Pearl (1997) to be
$\max\begin{Bmatrix}p_{10|1}\\ p_{10|0}\\ p_{10|0}+p_{11|0}-p_{00|1}-p_{11|1}\\ p_{01|0}+p_{10|0}-p_{00|1}-p_{01|1}\end{Bmatrix}\leq Q\{Y_{0}=1\}\leq\min\begin{Bmatrix}1-p_{00|1}\\ 1-p_{00|0}\\ p_{01|0}+p_{10|0}+p_{10|1}+p_{11|1}\\ p_{10|0}+p_{11|0}+p_{01|1}+p_{10|1}\end{Bmatrix}$ (31)
and
$\max\begin{Bmatrix}p_{11|0}\\ p_{11|1}\\ -p_{00|0}-p_{01|0}+p_{00|1}+p_{11|1}\\ -p_{01|0}-p_{10|0}+p_{10|1}+p_{11|1}\end{Bmatrix}\leq Q\{Y_{1}=1\}\leq\min\begin{Bmatrix}1-p_{01|1}\\ 1-p_{01|0}\\ p_{00|0}+p_{11|0}+p_{10|1}+p_{11|1}\\ p_{10|0}+p_{11|0}+p_{00|1}+p_{11|1}\end{Bmatrix}~{}.$ (32)
It follows from $\mathbf{Q}\subseteq\mathbf{Q}_{1}^{\ast}$ that $\Theta_{0}(P,\mathbf{Q})\subseteq\Theta_{0}(P,\mathbf{Q}_{1}^{\ast})$, and thus (30) is contained in (31)–(32).
Next, we show that there exists a $P$ for which $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$, $\Theta_{0}(P,\mathbf{Q})$ is not given by (5) and $\Theta_{0}(P,\mathbf{Q})\subsetneqq\Theta_{0}(P,\mathbf{Q}_{1}^{\ast})$. We do so by providing a numerical example. Consider the $P$ specified in Table 1 and the $Q$ specified in Table 2, where we write $q(y_{0}y_{1},d_{0}d_{1})=Q\{Y_{d}=y_{d},D_{z}=d_{z},~{}(d,z)\in\mathcal{D}\times\mathcal{Z}\}$ and omit any $q(\cdot)=0$. One can check that $Q\in\mathbf{Q}$ and rationalizes $P$, so that $Q\in\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$. In this example, $Q\{Y_{0}=1\}=0.4274$, and thus the identified set for $Q\{Y_{0}=1\}$ relative to $\mathbf{Q}$ is the singleton $\{0.4274\}$. In contrast, evaluating (31) at $P$ gives the identified set for $Q\{Y_{0}=1\}$ relative to $\mathbf{Q}_{1}^{*}$ as $[0.3336,0.5212]$. We thus conclude that $\Theta_{0}(P,\mathbf{Q})\subsetneqq\Theta_{0}(P,\mathbf{Q}_{1}^{\ast})$ for some $P$ that can be rationalized by $Q\in\mathbf{Q}$. Now consider evaluating the bounds of (5) at the same $P$. Doing so results in bounds on $Q\{Y_{0}=1\}$ given by $[0.1618,0.5445]$ if setting $z^{*}(0)=0$ and given by $[0.2656,0.8829]$ if setting $z^{*}(0)=1$. Therefore, no matter $z^{*}(0)=0$ or $1$, the bounds (5) is not the identified set for $Q\{Y_{0}=1\}$ relative to either $\mathbf{Q}_{1}^{*}$ or $\mathbf{Q}$.
$p_{00|0}$ $p_{10|0}$ $p_{01|0}$ $p_{11|0}$
0.4555 0.1618 0.3077 0.0750
$p_{00|1}$ $p_{10|1}$ $p_{01|1}$ $p_{11|1}$
0.1171 0.2656 0.0188 0.5985
Table 1: Distribution $P$ in Appendix B.2.
$q(00,01)$ $q(00,10)$ $q(01,01)$ $q(01,10)$
0.0039 0.0428 0.4516 0.0743
$q(10,01)$ $q(10,10)$ $q(11,01)$ $q(11,10)$
0.0149 0.2649 0.1469 0.0007
Table 2: Distribution $Q$.
B.3 Details of Example 5.2
Suppose $\mathcal{Y}=\{0,1\}$, $\mathcal{D}=\{0,1,2\}$, and $\mathcal{Z}=\{0,1,2\}$. Then, the linear program approach in Balke and Pearl (1993, 1997) leads to the following identified set for $\mathbb{E}_{Q}[Y_{1}]=Q\{Y_{1}=1\}$ relative to $\mathbf{Q}$ being defined as in Example 5.2:
$\left[\max\begin{Bmatrix}p_{11|0}\\ p_{11|1}\\ p_{11|2}\\ p_{11|0}-p_{11|1}+p_{11|2}\end{Bmatrix},\quad\min\begin{Bmatrix}1-p_{01|2}\\ 1-p_{01|1}\\ 1-p_{01|0}\\ 1-p_{01|0}+p_{01|1}-p_{01|2}\end{Bmatrix}\right]~{}.$ (33)
We will show that for some $P$ such that $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$, we have $\Theta_{0}(P,\mathbf{Q})$ strictly smaller than (5) and $\Theta_{0}(P,\mathbf{Q}^{*}_{1})$. For this purpose we are only concerned with the validity of (33) instead of its sharpness. For the lower bounds, first note for $z\in\mathcal{Z}$,
$Q\{Y_{1}=1\}=Q\{Y_{1}=1|Z=z\}\geq Q\{Y_{1}=1,D=1|Z=z\}=Q\{Y=1,D=1|Z=z\}~{},$
and therefore the first three rows follow. To show the last row, note it’s equivalent to
$Q\{Y_{1}=1,D_{1}=1\}+Q\{Y_{1}=1,D_{0}=0\}+Q\{Y_{1}=1,D_{0}=2\}\geq Q\{Y_{1}=1,D_{2}=1\}~{}.$
It therefore suffices to show that
$\{D_{2}=1\}\implies\{D_{1}=1\}\cup\{D_{0}=0\}\cup\{D_{0}=2\}~{}.$ (34)
Suppose $D_{2}=1$ but $D_{0}\neq 0$ and $D_{0}\neq 2$. Then $D_{0}=1$. But $D_{0}\leq D_{1}\leq D_{2}$, so $D_{1}=1$. (34) now follows. The lower bounds in (33) have all been shown to hold, and the upper bounds can be proved similarly.
[…]
$p_{00|0}$ $p_{10|0}$ $p_{01|0}$ $p_{11|0}$ $p_{02|0}$ $p_{12|0}$
0.3808 0.2427 0.0745 0.1664 0.0345 0.1011
$p_{00|1}$ $p_{10|1}$ $p_{01|1}$ $p_{11|1}$ $p_{02|1}$ $p_{12|1}$
0.2830 0.1947 0.0689 0.0712 0.2014 0.1808
$p_{00|2}$ $p_{10|2}$ $p_{01|2}$ $p_{11|2}$ $p_{02|2}$ $p_{12|2}$
0.0802 0.0976 0.1739 0.1165 0.2444 0.2874
Table 3: Distribution $P$ in Appendix B.3.
$q(000,002)$ $q(000,011)$ $q(000,021)$ $q(001,002)$ $q(001,020)$ $q(001,021)$ $q(001,121)$ $q(001,211)$
0.0139 0.0304 0.0054 0.2238 0.0802 0.0271 0.0735 0.0375
$q(010,101)$ $q(010,111)$ $q(010,122)$ $q(100,000)$ $q(100,022)$ $q(100,110)$ $q(100,202)$ $q(101,202)$
0.0453 0.0712 0.0499 0.0966 0.1461 0.0010 0.0345 0.0636
Table 4: Distribution $Q_{\text{ex},\text{min}}$.
$q(001,021)$ $q(001,121)$ $q(001,211)$ $q(010,000)$ $q(010,001)$ $q(010,002)$ $q(010,010)$ $q(010,122)$
0.0305 0.0745 0.0689 0.0220 0.0064 0.0085 0.0260 0.1664
$q(010,202)$ $q(011,002)$ $q(011,022)$ $q(011,210)$ $q(110,000)$ $q(110,001)$ $q(110,011)$ $q(110,022)$
0.0345 0.2116 0.0758 0.0322 0.0976 0.0971 0.0130 0.0350
Table 5: Distribution $Q_{\text{ex},\text{max}}$.
$q(000,000)$ $q(000,001)$ $q(000,002)$ $q(000,022)$ $q(000,111)$ $q(000,122)$ $q(000,222)$ $q(001,002)$
0.0802 0.0079 0.0430 0.0181 0.0209 0.0536 0.0345 0.1066
$q(001,022)$ $q(001,222)$ $q(010,001)$ $q(010,111)$ $q(010,122)$ $q(100,000)$ $q(100,001)$ $q(100,011)$
0.0797 0.1011 0.0453 0.0712 0.0952 0.0976 0.0971 0.0480
Table 6: Distribution $Q_{\text{ex},\text{om},\text{min}}$.
$q(000,001)$ $q(000,111)$ $q(000,122)$ $q(010,000)$ $q(010,001)$ $q(010,012)$ $q(010,111)$ $q(010,122)$
0.0079 0.0689 0.0056 0.0802 0.1114 0.0430 0.0051 0.1613
$q(010,222)$ $q(011,002)$ $q(011,022)$ $q(011,222)$ $q(100,001)$ $q(110,000)$ $q(111,012)$ $q(111,022)$
0.0345 0.0835 0.0548 0.1011 0.0971 0.0976 0.0231 0.0249
Table 7: Distribution $Q_{\text{ex},\text{om},\text{max}}$.
B.4 Details of Example 5.3
Consider the ARUM defined by (10) with strict one-to-one targeting and $|\mathcal{D}|=3$, $|\mathcal{Z}|=|\mathcal{D}^{\dagger}|=2$. In this case, there are two targeted treatments and one non-targeted treatment. Following Lee and Salanié (2023), label that non-targeted treatment as treatment 00 and impose the normalization that $g(z,0)=0$ for all $z\in\mathcal{Z}$. Label $Z=0$ as the instrument value that targets treatment $1$ and label $Z=1$ as the instrument value that targets treatment $2$, so that (25) holds for $d=1,2$ for some $\overline{U}(d),\underline{U}(d)$ with $\overline{U}(d)>\underline{U}(d)$ and with $z^{\dagger}(1)=0$, $z^{\dagger}(2)=1$. Let $\mathbf{Q}$ denote the set of all distributions for which $(D_{z}:z\in\mathcal{Z})$ is determined by (10) with these restrictions and additionally imposing that the support of $(U_{0},U_{1},U_{2})$ is $\Re^{3}$. Recall $(U_{0},U_{1},U_{2})\perp\!\!\!\perp Z$ by assumption.
Let $U_{10}=U_{1}-U_{0}$ and $U_{20}=U_{2}-U_{0}$. In this model, the treatment value is completely determined by the vector of realizations $(U_{10},U_{20})$. For instance, $D_{z}=2$ if and only if
$U_{20}$ $\geq-g(z,2)$
$U_{20}-U_{10}$ $\geq g(z,1)-g(z,2)~{},$
and a similar characterization holds for $D_{z}=1$. See Figure 1, which is taken from Figure 1 in Lee and Salanié (2023).
[…]
Figure 1: Treatment under each value of $(u_{10},u_{20})$ for a given $z$.
We first show that (3) holds for the targeted treatments by verifying that (11) holds for the targeted treatments. Consider (11) for $d=1$. It holds with $z^{\ast}(1)=z^{\dagger}(1)=0$ because
$\overline{U}(1)>$ $~{}\underline{U}(1)~{},$
$\overline{U}(1)-\underline{U}(2)>$ $~{}\underline{U}(1)-\overline{U}(2)~{},$
which in turn holds because $\overline{U}(d)>\underline{U}(d)$ for $d\in\{1,2\}$. Thus (11) holds for $d=1$, which, as shown in Example 4.5, implies that (3) holds for $d=1$. By a parallel argument, (3) holds for $d=2$.
We now show that there does not exist a value of $z^{\ast}(0)$ such that (3) holds for the non-targeted treatment, treatment 00. Suppose $z^{\ast}(0)=0$. Then
$Q\{D_{0}\neq 0,~{}D_{1}=0\}\geq$ $~{}Q\{D_{0}=1,~{}D_{1}=0\}$
$=$ $~{}Q\{-\underline{U}(1)\geq U_{10}\geq-\overline{U}(1),U_{20}\leq-\overline{U}(2),U_{10}-U_{20}\geq\underline{U}(2)-\overline{U}(1)\}$
$>$ $~{}0~{},$
where the last line is using that the support of the distribution of $(U_{10},U_{20})=\Re^{2}$ by assumption and that strict targeting of treatment $1$ requires $-\underline{U}(1)>-\overline{U}(1)$. Thus (3) cannot hold for $d=0$ with $z^{\ast}(0)=0$. A parallel argument shows that (3) cannot hold for $d=0$ with $z^{\ast}(0)=1$.
We conclude that, when $|\mathcal{D}|=3$ and $|\mathcal{D}^{\dagger}|=|\mathcal{Z}|=2$, one-to-one strict targeting with the regularity condition that the support of $(U_{0},U_{1},U_{2})$ is $\Re^{3}$ implies that (3) holds for the targeted treatments but not for the non-targeted treatments, and thus Assumption 2.2 cannot hold. This argument can be adapted for any ARUM with $|\mathcal{D}|\geq 3$ and $|\mathcal{Z}|=|\mathcal{D}^{\dagger}|$ to show that, while (3) holds for the targeted treatments, there does not exist a value of $z^{\ast}(d)$ such that (3) holds for any non-targeted treatment $d$, and thus that Assumption 2.2 cannot hold.
$z=0$$z=1$$(1,1)$$(2,2)$$(0,0)$$(1,2)$$(0,2)$$(1,0)$
Figure 2: Values of $(D_{0},D_{1})$ for each value of $(u_{10},u_{20})$.
Next, consider the identified sets for the average potential outcomes. Since (3) is satisfied for the targeted treatments, a straightforward modification of the arguments underlying Theorem 3.2 show that the identified sets for $\mathbb{E}_{Q}[Y_{d}]$ for $d\in\{1,2\}$ is given by (5) for any $P$ such that $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$.
We now derive the identified set for the average potential outcome of the non-targeted treatment. First, note that $-g(0,1)=-\overline{U}(1)<-\underline{U}(1)=-g(1,1)$ and $-g(0,2)=-\underline{U}(2)>-\overline{U}(2)=-g(1,2)$. Therefore, it can be verified from Figure 2 that for all $Q\in\mathbf{Q}$,
$Q\{(D_{0},D_{1})\in\{(0,0),(1,0),(1,1),(0,2),(1,2),(2,2)\}\}=1~{}.$ (35)
Let $\mathbf{Q}^{\prime}$ denote the set of all distributions that satisfies (35). Note that all $Q\in\mathbf{Q}$ satisfies (35), so $\mathbf{Q}\subseteq\mathbf{Q}^{\prime}$. On the other hand, by assigning appropriate probabilities to each set in the partition in Figure 2, we immediately see that each $Q\in\mathbf{Q}^{\prime}$ can be rationalized by a $Q\in\mathbf{Q}$. Therefore, $\mathbf{Q}=\mathbf{Q}^{\prime}$. Using linear programming as in Balke and Pearl (1993, 1997), we obtain the following identified set for $\mathbb{E}_{Q}[Y_{0}]=Q\{Y_{0}=1\}$ relative to $\mathbf{Q}$:
$\left[\max\begin{Bmatrix}p_{10|0}\\ p_{10|1}\end{Bmatrix},\quad\min\begin{Bmatrix}1-p_{00|1}\\ 1-p_{00|0}\end{Bmatrix}\right]~{}.$ (36)
The identified set in (36) equals (7) for $d=0$ with $Y$ and $Z$ binary. Thus, the identified set for $\mathbb{E}_{Q}[Y_{0}]$ relative to $\mathbf{Q}$ corresponds to the identified set relative to $\mathbf{Q}^{\ast}_{3}$, the set of distributions that satisfy mean independence, 3.1. By the same sandwich argument used to prove Theorem 3.3, the identified set for $\mathbb{E}_{Q}[Y_{0}]$ relative to $\mathbf{Q}$ corresponds to the identified set relative to $\mathbf{Q}^{\ast}_{1}$, and thus imposing this ARUM has no identifying power for $\mathbb{E}_{Q}[Y_{0}]$ beyond instrument exogeneity.
Finally, we show that there exists a $P$ for which $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$ and (36) is strictly smaller than (5), so that $\Theta_{0}(P,\mathbf{Q})$ is not given by (5). We do so by providing a numerical example. Consider the $P$ specified in Table 8 and the $Q_{\text{arum},\text{min}}$ and $Q_{\text{arum},\text{max}}$ specified in Tables 9 and 10 respectively, where we write $q(y_{0}y_{1}y_{2},d_{0}d_{1})=Q\{Y_{d}=y_{d},D_{z}=d_{z},~{}(d,z)\in\mathcal{D}\times\mathcal{Z}\}$ and omit any $q(\cdot)=0$. One can check that both $Q_{\text{arum},\text{min}}$ and $Q_{\text{arum},\text{max}}$ rationalize $P$ and satisfy Assumption 2.1. One can further check that both $Q_{\text{arum},\text{min}}$ and $Q_{\text{arum},\text{max}}$ satisfy the restriction in (35), so that $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$. Evaluating (36) at $P$ gives the identified set for $\mathbb{E}_{Q}[Y_{0}]$ relative to $\mathbf{Q}$ as $[0.2518,0.8167]$. One can further check that the two endpoints are attained by $\mathbb{E}_{Q_{\text{arum},\text{min}}}[Y_{0}]=0.2518$ and $\mathbb{E}_{Q_{\text{arum},\text{max}}}[Y_{0}]=0.8167$. On the other hand, if one evaluates (5) by setting $z^{*}(0)=0,1$ at the same $P$, the resulting bounds for $\mathbb{E}_{Q}[Y_{0}]$ equal $[p_{10|0},1-p_{00|0}]=[0.2518,0.8937]$ and $[p_{10|1},1-p_{00|1}]=[0.2372,0.8167]$ respectively. In both cases, (36) is strictly contained in (5). Hence, $\Theta_{0}(P,\mathbf{Q})$ is not given by (5).
$p_{00|0}$ $p_{10|0}$ $p_{01|0}$ $p_{11|0}$ $p_{02|0}$ $p_{12|0}$
0.1063 0.2518 0.2946 0.3183 0.0020 0.0270
$p_{00|1}$ $p_{10|1}$ $p_{01|1}$ $p_{11|1}$ $p_{02|1}$ $p_{12|1}$
0.1833 0.2372 0.0140 0.1399 0.1701 0.2555
Table 8: Distribution $P$ in Appendix B.4.
$q(000,10)$ $q(000,11)$ $q(000,12)$ $q(000,22)$ $q(001,02)$ $q(001,12)$
0.0049 0.0140 0.1535 0.0020 0.1063 0.1222
$q(001,22)$ $q(010,10)$ $q(010,11)$ $q(100,00)$ $q(100,02)$
0.0270 0.1784 0.1399 0.2372 0.0146
Table 9: Distribution $Q_{\text{arum},\text{min}}$.
$q(000,00)$ $q(010,10)$ $q(100,00)$ $q(100,10)$ $q(100,11)$ $q(100,02)$
0.1063 0.0770 0.0837 0.0521 0.0140 0.1681
$q(100,22)$ $q(101,12)$ $q(101,22)$ $q(110,10)$ $q(110,11)$
0.0020 0.2285 0.0270 0.1014 0.1399
Table 10: Distribution $Q_{\text{arum},\text{max}}$.